### v3

# Spatial-RAG: Spatial Retrieval Augmented Generation for Real-World Spatial Reasoning Questions

Abstract
Spatial reasoning remains challenging for Large Language Models (LLMs), which struggle with spatial data retrieval and reasoning. We propose Spatial Retrieval-Augmented Generation (Spatial-RAG), a framework that extends RAG to spatial tasks by integrating sparse spatial retrieval (spatial databases) and dense semantic retrieval (LLM-based similarity). A multi-objective ranking strategy balances spatial constraints and semantic relevance, while an LLM-guided generator ensures coherent responses. Experiments on a real-world tourism dataset show that Spatial-RAG significantly improves spatial question answering, bridging the gap between LLMs and spatial intelligence.
RAG, spatial, LLM agent, tourism question answering
1 Introduction
Spatial questioning has long been a fundamental domain (Obe & Hsu, 2021), which adeptly handles a variety of spatial questions, from identifying closest neighbors to detecting line-polygon intersections. However, traditional spatial questioning systems rely on specialized spatial query languages, which are vastly different from human language, making them inaccessible to layman users. More critically, these systems lack the ability to infer complex spatial and semantic relationships from the nuanced, context-rich nature of human text, limiting their applicability to real-world question-answering scenarios. Recent advancements in Large Language Models (LLMs) have transformed many fields within machine learning (ML), particularly in understanding and generating human-like text. This progress has inspired early efforts to bridge the gap between spatial questioning and natural language by directly eliciting spatial knowledge from LLMs. These efforts span a range of applications, including geographic encyclopedic question answering (Mai et al., 2021, 2020; Chen, 2014; Scheider et al., 2021; Contractor et al., 2021a), geolocalization (Zhou et al., 2024; Haas et al., 2024), and automated High-definition map generation (Li et al., 2022; Liu et al., 2023). Despite these advancements, recent studies indicate that LLMs exhibit significant limitations in spatial reasoning (Mai et al., 2024; Roberts et al., 2023), and struggle with even basic spatial tasks, such as geoparsing (Mai et al., 2024) and understanding relative spatial relationships (Majic et al., 2024). This gap becomes particularly evident when handling real-world spatial reasoning tasks, such as the one illustrated in Figure 1, where the system must recommend a restaurant along a specific route within a predefined area.
Figure 1: An example of a real-world spatial reasoning question with nearby spatial objects. Areas that satisfy the spatial constraint are highlighted in blue.
This paper tackles a novel challenge: augmenting LLMs with spatial reasoning capabilities. Specifically, we extend Retrieval-Augmented Generation (RAG) into spatial information retrieval and reasoning, bridging the gap between structured spatial databases and unstructured textual reasoning. RAG has demonstrated its effectiveness in knowledge-intensive tasks, such as question answering (QA) (Siriwardhana et al., 2023), by retrieving domain-specific documents to enhance LLM responses. However, existing RAG systems primarily focus on retrieving and generating textual content and lack the spatial intelligence required for spatial reasoning tasks, especially tasks that involve understanding and computing complex spatial relationships among geometries, including points, polylines, and polygons.
Spatial questioning is intrinsically challenging, requiring a synergistic combination of text-based inference and spatial computations. For example, as illustrated in Figure 1, answering the question requires LLM to elicit and formulate the user’s textual request into the problem of ”finding points near the polyline” and solve it based on a spatial map (database). Then, it also requires inferring user preference to select the spatially and semantically preferred candidates. Thus, the system must seamlessly integrate structured spatial retrieval with unstructured text-based reasoning, ensuring both spatial accuracy and contextual understanding. To achieve this, we introduce Spatial Retrieval-Augmented Generation (Spatial-RAG), a novel framework that unifies text-guided spatial retrieval with spatially aware text generation. Specifically, to identify spatially relevant candidate answers, we propose a novel spatial hybrid retrieval module combining sparse and dense retrievers. To rank the candidates and generate the final answers, we propose to fuel the generator with retrieved results on the Pareto frontier based on a spatial and semantic joint ranking strategy. Our contributions are summarized as follows:
  • A generic Spatial RAG framework: We introduce spatial-RAG, the first framework that extends RAG to spatial question answering, to tackle a broad spectrum of spatial reasoning tasks, such as geographic recommendation, spatially constrained search, and contextual route planning. Our approach seamlessly integrates spatial databases, LLMs, and retrieval-based augmentation, enabling effective handling of complex spatial reasoning questions directly within the familiar operational paradigm of LLMs.
  • Sparse-dense spatial hybrid retriever: We propose a hybrid retrieval mechanism that combines sparse retrieval (SQL-based structured queries) with dense retrieval (LLM-powered semantic matching). This dual approach ensures that retrieved results align both spatially and semantically with the user’s query, significantly improving retrieval accuracy in spatial contexts.
  • Multi-objective guided spatial text generator: To handle both spatial constraints and textual reasoning in the spatial question-answering task, we introduce a multi-objective optimization framework that dynamically balances trade-offs between spatial and semantic relevance. This ensures that generated responses are both geometrically accurate and linguistically coherent.
  • Real-World Evaluation: We evaluated our method on a real-world dataset collected from a Tourism website that has user questions and reviews about different spatial entities. Experiments on this dataset reveal the ability to handle real-world spatial reasoning questions.
Through these innovations, Spatial-RAG significantly enhances the spatial reasoning capabilities of LLMs, bridging the gap between structured spatial databases and natural language question answering.
2 Related Work
2.1 Retrieval Augmented Generation
Retrieval-Augmented Generation (RAG) is a hybrid approach that integrates retrieval systems and generative models to enhance factual accuracy and contextual relevance in natural language generation (Fan et al., 2024). Unlike conventional language models that rely solely on parametric memory, RAG dynamically retrieves relevant external knowledge before generating a response. One of the foundational works in RAG is Lewis et al. (2020), where a retrieval module fetches relevant passages from a large-scale knowledge corpus (e.g., Wikipedia), which are then fused with the question context to generate a more informed response. This technique has proven particularly effective in open-domain question answering (QA), fact verification, and context-aware text generation. RAG systems have expanded beyond text and document retrieval to incorporate a wide variety of data types (He et al., 2024) — tables, graphs, charts, and diagrams. While RAG has been widely explored, its application in spatial reasoning question answering remains an unexplored research area. Existing studies have primarily focused on knowledge-grounded dialogues (Yu et al., 2024) but often struggle with integrating spatial computation into the question-answering process effectively.
2.2 Spatial Questions
Spatial questions in domain-specific applications can generally be categorized into two distinct types: 1) Textual Knowledge-based Spatial Questions These are spatial questions that can be answered by traditional QA methods without the need for spatial computation and reasoning (Liétard et al., 2021). For example, the question ”What is the population of Los Angeles city?” falls under this category. Despite their spatial context, these questions are essentially text-based and, hence, can be effectively addressed using traditional Retrieval-Augmented Generation (RAG) methods (Christmann & Weikum, 2024). 2) Spatial Reasoning Questions This category encapsulates spatial questions that demand a model’s capability to comprehend and reason with spatial data and spatial relationships. A common example is a model being presented with textual information describing the spatial relationships among multiple objects (Li et al., 2024). An example question could be, ”What is the position of object A relative to object B?”, where objects A and B are locations or entities specified on the map. Resolving such queries requires a profound understanding of spatial concepts and robust reasoning skills, which largely depend on the model’s training to handle spatial data. Several studies (Mai et al., 2024; Roberts et al., 2023) have investigated the capacity of LLMs to understand spatial concepts, yet these models often struggle with accurate reasoning even after fine-tuning. Other research (Li et al., 2023) has attempted to enhance this ability by converting geolocation coordinates into addresses to enrich the semantic context. However, these improvements tend to be marginal and are mostly limited to straightforward reasoning tasks like describing positions. Moreover, many existing methods rely on predefined sets of actions tailored to specific tasks.
3 Problem Formulation
In this study, our primary focus is Spatial Reasoning Questions. We formulate the problem as follows: Given a question $q$, the system aims to generate an answer $y$,
$y^{*}$ $=\arg\max_{y}\ \lambda_{s}^{T}f_{s}(q,y)+\lambda_{k}^{T}f_{k}(q,y)$ (1)
s.t. $y\in C_{s}(q),\quad y\in C_{k}(q),$
$\lambda_{s}\geq 0,\quad\lambda_{k}\geq 0,$
$\mathbf{1}^{T}\lambda_{s}+\mathbf{1}^{T}\lambda_{k}=1,$
where $f_{s}\in\mathbb{R}^{d_{s}}$ is the spatial relevance score vector, $f_{k}\in\mathbb{R}^{d_{k}}$ is the semantic relevance score vector, $C_{s}$ is the spatial candidate set that satisfies the spatial constraints of the question, $C_{k}$ is the semantic candidate set that satisfies the semantic constraints of the question, $\lambda_{s},\lambda_{k}$ are the spatial weights and semantic weights, respectively, $y^{*}$ is the optimal answer, $\mathbf{1}^{T}\lambda_{s}+\mathbf{1}^{T}\lambda_{k}=1$ ensures a normalized trade-off.
To solve this problem, there are three questions to answer: 1) How to acquire the spatial candidate set $C_{s}$? 2) How to synergize the spatial and texts by evaluating $f_{s}(q,y)$? 3) How to trade off the spatial and semantic aspects?
4 Methodology
Figure 2: Illustration of the proposed Spatial-RAG framework.
4.1 Overview
Our proposed framework, Spatial Retrieval-Augmented Generation (Spatial-RAG), is illustrated in Figure 2. Spatial-RAG consists of three key stages: First, to construct the spatial candidate set $C_{s}$, the system must precisely define spatial constraints and then retrieve spatial objects that satisfy them. As depicted in Figure 2 (Sparse Spatial Retrieval), we achieve this by parsing the input natural language questions into a spatial SQL query, which will be executed on the spatial database to efficiently retrieve relevant spatial objects from the database. This process is detailed in Section 4.2. Second, to effectively compute spatial relevance $f_{s}(q,y)$ while integrating textual information, we propose a hybrid spatial retrieval scheme. As shown in Figure 2 (Dense Retrieval Module), this method combines sparse spatial relevance scores from the database with dense semantic similarity scores from text embeddings. This enables the system to rank retrieved spatial objects based on their spatial relevance to the input question, as detailed in Section 4.3. Third, given both spatial and semantic constraints, we formulate a multi-objective optimization problem to balance these factors. The system computes the Pareto front of candidate answers, and the LLM dynamically trades off among these solutions to generate an optimal response. This step is covered in Section 4.4.
4.2 Sparse Spatial Retrieval
The answer to a spatial reasoning question must meet specific spatial constraints. The spatial candidate set $C_{s}(q)$ consists of all possible answers $y$ that satisfy a set of spatial constraints $\mathcal{C}_{s}(q)$. Formally, we define:
$C_{s}(q)=\{y\mid{c}_{s}(y,q)\leq 0,\forall{c}_{s}\in\mathcal{C}_{s}(q)\},$ (2)
where ${c}_{s}(y,q)$ represents a constraint function that encodes a spatial condition (e.g., topological, directional, or distance-based constraints), $\mathcal{C}_{s}(q)$ is the set of all spatial constraints associated with the question $q$. For example, if the spatial constraint requires $y$ to be within a distance $\epsilon$ from a reference location $l_{q}$, then a possible constraint function is:
$c_{s}(y,q)=\text{d}(y,l_{q})-\epsilon\leq 0.$ (3)
This formulation ensures that only spatially valid answers are included in $C_{s}(q)$.
Addressing spatial constraints requires executing a well-defined spatial SQL query within a spatial database. This process involves identifying the appropriate query function, the reference spatial objects, the target spatial objects, and any necessary numerical parameters. Formally, a spatial SQL query can be expressed as:
$Q_{s}=\mathcal{F}_{s}(G_{r},G_{t},\epsilon)$ (4)
where $\mathcal{F}_{s}$ is the spatial query function that determines the relationship between objects. $G_{r}$ represents the set of reference objects extracted from the question. $G_{t}$ represents the set of target objects that are potential answers. $\epsilon$ is the set of numerical parameters and spatial relationships governing the spatial constraint (e.g., distance threshold, topological relations).
Given the diversity and potentially complex nature of these constraints, Large Language Models (LLMs) often struggle to directly construct a complete and executable spatial query from user input. To bridge this gap, we structure the spatial query incrementally, allowing the LLM to systematically populate the required components.
Our approach follows three key steps: 1) Geometry Recognition: Identify and extract the reference spatial objects $G_{r}$ and candidate target spatial objects $G_{t}$ from the user’s input and extract their spatial footprints – geometries. 2) Query Function Selection: Determine the appropriate spatial function $\mathcal{F}_{s}$ based on the intended spatial relationship (e.g., containment, proximity). 3) Parameter Estimation: Assign numerical constraints $\epsilon$ to ensure precise spatial filtering (e.g., buffer radius).
By formalizing this structured process, we enhance the LLM’s ability to generate accurate and executable spatial SQL queries. This, in turn, improves the system’s capability to handle complex spatial reasoning questions effectively.
4.2.1 Geometry Recognition
In spatial reasoning tasks, accurately identifying spatial objects and extracting their spatial footprints (i.e., geometries) are essential for parsing questions to spatial queries. Spatial footprints of spatial objects, denoted as $g\in\mathcal{G}$, can generally be categorized into three fundamental types: points, polylines, and polygons. Formally, we define these categories as follows:
  • Point: $\mathcal{G}_{\text{point}}=\{g\mid g\in\mathbb{R}^{2},\dim(g)=0\}$ This category includes single points and multipoints, representing locations with negligible area. Examples include stop signs, address points, and a user’s current location. In spatial databases, these entities are typically represented as the ’Point’ geometry type.
  • Polyline: $\mathcal{G}_{\text{line}}=\{g\mid g\subseteq\mathbb{R}^{2},\dim(g)=1\}$ Polylines, including multipolylines, represent linear one-dimensional objects with negligible width. Common examples include streets, streams, bus routes, and power lines. In spatial databases, these geometries are abstracted as the ’LineString’ type.
  • Polygon: $\mathcal{G}_{\text{polygon}}=\{g\mid g\subseteq\mathbb{R}^{2},\dim(g)=2\}$ Polygons, including multipolygons, represent two-dimensional objects that define enclosed areas. These geometries are essential for depicting regions such as census areas, parcels, counties, neighborhoods, and zoning areas.
The complexity of a spatial query depends on the types of spatial footprints of objects involved. For simpler queries, such as ”finding the nearest bus stop from a given location”, only point geometries are required, and the spatial candidate set is
$C_{s}=\{g|g\in\mathcal{G}_{\text{point}},d(g,g_{\text{point}})<\epsilon\}$ (5)
where $g_{\text{point}}\subseteq\mathcal{G}_{\text{point}}$ represents a point object (e.g., given location), $\epsilon$ is the distance threshold. For more complex queries, such as ”I will walk from home to the university campus along 7th Street and Jones Street; please recommend a café where I can buy breakfast on my walk.”, multiple geometry types must be considered, and the spatial candidate set is
$C_{s}=\{g|g\in\mathcal{G}_{\text{point}},g\in B(g_{\text{polyline}},\epsilon)\cup g_{\text{polygon}}\}$ (6)
where $g_{\text{polyline}}\subseteq\mathcal{G}_{\text{polygon}}$ represents a polyline object (e.g., a route), $g_{\text{polygon}}\subseteq\mathcal{G}_{\text{polygon}}$ represents a polygonal region (e.g., a university campus), $B$ is a buffer around the polyline, $\epsilon$ is the buffer size.
By structuring spatial queries in this way, we ensure precise geometric representation, facilitating robust spatial reasoning and query execution.
4.2.2 Query Function Recognition and Parameter Estimation
After recognizing the geometries involved in a spatial query, the subsequent step is to determine the appropriate spatial query functions $\mathcal{F}_{s}$ required to handle various geometrical interactions. Despite the differing interactions among geometries, these can be uniformly addressed using distance functions $\text{d}(g_{r},g_{t})$, which calculate the shortest distance between two geometrical entities $g_{r},g_{t}\in\mathcal{G}$.
Formally, given sets of reference geometries $G_{r}\subseteq\mathcal{G}$ and target geometries $G_{t}\subseteq\mathcal{G}$ , the spatial candidate set $C_{s}$ can be defined as:
$\begin{cases}\{g_{t}\in G_{t}\mid\exists g_{r}\in G_{r},\text{d}(g_{r},g_{t})\leq\epsilon\},&\text{if }\text{d}(g_{r},g_{t})>0,\\ \{g_{t}\in G_{t}\mid\exists g_{r}\in G_{r},g_{r}\cap g_{t}\neq\emptyset\},&\text{if }\text{d}(g_{r},g_{t})=0.\end{cases}$ (7)
Parameters such as search radius or buffer distance $\epsilon$ are autonomously determined by the LLM, typically grounded in contextual understanding (e.g., estimated walking distance or area of interest). The parameter $\epsilon$ can be represented as: $\epsilon=\phi(q)$, where $\phi$ is a function that maps the context of the query $q$ to an appropriate numerical value.
Once the geometries $G_{r}$, $G_{t}$, functions $\mathcal{F}_{s}$, and parameters $\epsilon$ are delineated, the system constructs the precise spatial query $Q_{s}$. This query can be formally expressed by Equation 4, which ensures exact retrievals from the spatial database, maintaining both accuracy and relevance in the results. By leveraging these mathematical formulations, the system effectively translates spatial reasoning tasks into executable queries, facilitating robust spatial intelligence within the LLM framework.
4.3 Hybrid Spatial Objects Ranking
The spatial relevance score $f_{s}$ consists of two components: a score derived from sparse spatial retrieval from the spatial database and a score from spatial dense retrieval based on text similarity between the question and the spatial descriptions of candidate objects. Formally, we define:
$f_{s}=\lambda_{s}f_{s}^{\text{sparse}}+\lambda_{d}f_{s}^{\text{dense}},$ (8)
where $\lambda_{s}$ and $\lambda_{d}$ are weighting coefficients controlling the contribution of each score.
4.3.1 Sparse Spatial Relevance Scoring
Sparse spatial relevance is computed directly from the spatial database using explicit spatial relationships. The score is determined by the spatial query function $\mathcal{F}_{s}$, which computes the distance between reference and target objects. Formally, we define:
$f_{s}^{\text{sparse}}=\begin{cases}\displaystyle\frac{1}{1+d(g_{r},g_{t})},&\text{if }g_{r}\cap g_{t}=\emptyset\\ 1,&\text{if }g_{r}\cap g_{t}\neq\emptyset\end{cases}$ (9)
where $g_{r}$ and $g_{t}$ are reference and target spatial objects, respectively. $d(g_{r},g_{t})$ is a distance function measuring proximity in the spatial database. If $g_{t}$ overlaps with $g_{r}$, we assign a perfect relevance score of 1.
This ensures that objects within a region are maximally relevant, while those outside the region receive scores that decay with increasing distance.
4.3.2 Dense Spatial Relevance Scoring
Unlike sparse scoring, dense spatial relevance is inferred from textual descriptions associated with spatial objects. We leverage an LLM to extract key spatial attributes from user queries and compare them with the descriptions of candidate objects.
Extracting Spatial Requirements
Given a user query $q$ and a set of text descriptions $d_{t}$ for spatial objects $G_{t}$, we extract the relevant spatial content via an attention-based masking function:
$v_{q,s}=\mathcal{E}(\mathcal{M}_{s}(q)),\quad v_{t,s}=\mathcal{E}(\mathcal{M}_{s}(d_{t})),$ (10)
where $v_{q,s}$ and $v_{t,s}$ are dense vector representations of spatial features, and $\mathcal{M}$ is the extraction function mapping input text to a spatial related text, $\mathcal{E}$ is the text encoder.
Ranking via Cosine Similarity
The relevance score is computed via cosine similarity:
$f_{s}^{\text{dense}}=\frac{v_{q,s}\cdot v_{t,s}}{\parallel v_{q,s}\parallel\parallel v_{t,s}\parallel}.$ (11)
4.3.3 Hybrid Ranking as a Generalized Model
We can demonstrate that hybrid ranking generalizes both sparse and dense ranking approaches:
  • Sparse-Only Case: If $\lambda_{d}=0$ , then $f_{s}=\lambda_{s}f_{s}^{\text{sparse}}$, reducing to a purely distance-based ranking.
  • Dense-Only Case: If $\lambda_{s}=0$ , then $f_{s}=\lambda_{d}f_{s}^{\text{dense}}$, reducing to a purely semantic-based ranking.
  • Hybrid Case (General): If both weights are nonzero, hybrid ranking benefits from both explicit spatial constraints and implicit semantic relevance, leading to a more comprehensive ranking mechanism.
This formulation ensures that hybrid ranking outperforms any single-ranking approach by capturing both spatial proximity and semantic alignment.
4.4 Multi-objective Generation
The semantic candidate set $C_{k}$ and the semantic relevance score $f_{k}$ are calculated based on dense vector similarity, we put the details in appendix A. After all the scores and candidate sets are acquired, the problem becomes a multi-objective optimization problem since each perspective (spatial and semantic) contributes independently.
4.4.1 Pareto Front Computation
Given the spatial and semantic relevance scores, our goal is to identify the Pareto-optimal candidates that achieve the best trade-off between these objectives. A candidate $y$ is Pareto-optimal if no other candidate dominates it in both spatial and semantic relevance. Formally, the Pareto front $P(q)$ is defined as:
[…]
$f_{s}(q,y^{\prime})\geq f_{s}(q,y)\ \text{and}\ f_{k}(q,y^{\prime})\geq f_{k}(q,y),$
[…]
This ensures that each candidate in $P(q)$ is non-dominated, meaning no other candidate is strictly better in both spatial and semantic relevance.
4.4.2 LLM-Based Trade-Off Decision
Once the Pareto front $P(q)$ is determined, we use an LLM to dynamically balance the trade-offs between spatial constraints and semantic preferences based on the context of the user query. Specifically, the LLM receives the user query, sparse spatial relevance scores, and spatial object descriptions as input:
$I=\left\{q,\left(f_{s}^{\text{sparse}}(q,y)\ ,d_{y}\right)\ ,\forall y\in P(q)\right\}.$ (13)
A dynamic weighting function $\lambda_{s},\lambda_{k}=h(I)$ based on contextual information is extracted from the input, adjusting the importance of spatial vs. semantic relevance, where $h$ is a learned function capturing query-specific trade-offs.
The top-ranked candidate $y^{*}$ is selected by LLM:
$y^{*}=\arg\max_{y\in P(q)}\lambda_{s}^{T}f_{s}(q,y)+\lambda_{k}^{T}f_{k}(q,y),$ (14)
and the LLM generates a natural language response.
The system adapts to different query contexts instead of using a fixed weighting scheme. By structuring decision-making into discrete steps (candidate filtering $\rightarrow$ Pareto selection $\rightarrow$ trade-off balancing $\rightarrow$ response generation), the LLM avoids generating infeasible or illogical results. This structured approach maximizes accuracy and usability, ensuring that the system’s final response aligns closely with the original user intent.
5 Experiment
5.1 Experiment Setting
5.1.1 Datasets
TourismQA (Contractor et al., 2019) dataset contains user questions crawled from TripAdvisor posts from 50 cities around the world. Reviews of restaurants, attractions, and hotels for each city are crawled from travel forums and hotel booking websites. We select two two popular tourist cities, New York City and Miami, to evaluate the performance of different methods. For New York City, the original dataset contains information on 9,470 Points of Interest (POIs), with a total of 17,448 QA pairs. For Miami, the dataset includes 2,640 POIs and 133 QA pairs. Our preprocessing steps on the dataset include removing POIs with empty review information and eliminating duplicate QA pairs.
5.1.2 Evaluation Metrics
We assess the experimental results from multiple dimensions: 1) Delivery Rate: Following (Xie et al., 2024), this metric assesses whether the method can deliver a result successfully. 2) Spatial Sparse Pass Rate: This metric evaluates whether the parsed spatial query is correct. 3) Spatial Dense Pass Rate: This metric evaluates whether the answer satisfies the spatial-related semantic constraints in the question. 4) Semantic Pass Rate: This metric evaluates whether the answer meets the semantic constraints in the question. We use GPT-4o to evaluate the last three metrics.
5.1.3 Models for comparison
City Model Delivery Rate Spatial Sparse Pass Rate Spatial Dense Pass Rate Semantic Pass Rate
NYC SD 100 - 57.1 32.1
TE 99.6 - 48.6 54.6
ST 100 - 53.6 55.9
Naive RAG 99.8 - 52.0 54.5
GeoLLM 99.8 - 69.5 42.8
Spatial-RAG (GPT-3.5-Turbo) 87.2 67.0 64.4 47.4
Spatial-RAG (GPT-4-Turbo) 86.1 65.0 71.6 50.1
Miami SD 100 - 36.8 22.4
TE 100 - 28.9 39.5
ST 100 - 26.3 42.1
Naive RAG 100 - 31.6 40.8
GeoLLM 100 - 52.6 36.8
Spatial-RAG (GPT-3.5-Turbo) 86.8 75.8 57.6 43.9
Spatial-RAG (GPT-4-Turbo) 86.8 81.8 51.5 45.5
Table 1: Performance comparison of models in New York City (NYC) and Miami. The framework is deployed on GPT-4-Turbo and GPT-3.5-Turbo and compared its performance with several baseline models, including SD, TE, ST, Naive RAG, and GeoLLM.
Model Delivery Rate Spatial Sparse Pass Rate Spatial Dense Pass Rate Semantic Pass Rate
Spatial-RAG (GPT-4-Turbo) 86.1 65.0 71.6 50.1
w/o sparse spatial 98.9 - 53.3 78.4
w/o dense spatial 89.3 61.5 68.4 49.6
w/o dense semantic 85.9 72.8 75.9 34.8
Table 2: Different modules of the proposed framework are removed for ablation study.
We used GPT-3.5-Turbo and GPT-4-Turbo as the LLM in our framework to assess the impact of LLM capability on performance and compared it against the following four methods: Sort-by-distance (SD) (Contractor et al., 2021b): This method ranks the candidate spatial objects based on their distance to the reference objects in the spatial question. Text embedding (TE) (Cakaloglu et al., 2020): This greedy method minimizes the distance between the vector embeddings of the text description of the reference object and the target object. Spatial-text (ST): This approach computes the embeddings of the user’s question and compares the similarity between the question embedding and the text description embedding of the target object. Additionally, the object’s location is encoded as a distance score. The answer is then determined based on the weighted average of these scores. Naive RAG (Lewis et al., 2020): This method saves all spatial objects’ descriptions in a vector database and retrieves the most relevant objects based on vector similarity. GeoLLM (Manvi et al., 2024): This method encodes the spatial objects to address and enrich its context by adding spatial information of nearby spatial objects.
5.2 Main Results
The main results are summarized in 1 For the NYC dataset, Spatial-RAG demonstrates minimal differences in Delivery Rate when deployed with different LLMs. Approximately 86.1% of the questions are successfully processed. The failure cases where the framework is unable to generate valid recommendations can be classified into two main categories: Around 12.9% of the questions fail to retrieve any spatial objects from the spatial database, either due to polygon recognition errors or because the specified region in the SQL query lacks relevant spatial objects. The remaining 0.9% of cases stem from the LLM’s inability to correctly re-rank the retrieved results during the reranking process.
Additionally, the performance gap of our framework based on two LLMs in generating spatial SQL queries is insignificant. However, GPT-4-Turbo outperforms GPT-3.5-Turbo by approximately 7% in Spatial Dense Pass Rate and also demonstrates a slight advantage in Semantic Pass Rate, and both models perform slightly below the optimal values (by 5% and 8%, respectively). This is due to the trade-off made when jointly considering both spatial proximity and the user’s personal preferences.
Since the SD method returns the nearest spatial object, it achieves a higher spatial dense retrieval score than the TE, ST, and Naive RAG methods. However, its exclusive reliance on distance results in inferior performance on other metrics. For GeoLLM, as it considers only the names of spatial objects and their distances, it also achieves competitive performance in Spatial Dense Pass Rate. In contrast, both the TE and ST methods consider semantic contexts, leading to superior performance in this aspect compared to other models. The Naive RAG and ST methods jointly optimize for both spatial dense retrieval and user semantic dense retrieval, resulting in comparable performance between the two models.
For Miami, which differs significantly from NYC in terms of area and geography, Spatial-RAG also demonstrates strong performance. The performance of the baseline models almost remains consistent with the patterns observed in the previous analysis.
Figure 3: An example of how Spatial-RAG operates: Given a question, 1) Sparse Spatial Retrieval: an LLM parses the natural language question into a spatial SQL query for spatial database, retrieving spatial objects that meet the spatial constraints and the sparse spatial relevance scores. 2) Question Decomposition and Dense Retrieval: In parallel, Spatial-RAG decomposes the question into spatial and semantic components and compares them with the descriptions of spatial objects to perform dense retrieval, filtering out irrelevant ones. 3) LLM Reranking: A language agent balances the spatial and semantic aspects to rerank the candidate and generate the final answer.
5.3 Ablation Study
We conducted three ablation studies to assess the importance of different modules in our framework. We remove the sparse spatial module, dense spatial module, and dense semantic module, respectively, and denote the ablation versions as w/o sparse spatial, w/o dense spatial, and w/o dense semantic. The results are summarized in Table 2. It’s found that there is a significant increase in Delivery Rate after removing the sparse spatial module, as this eliminates the need to determine the geometry type. In this case, the model relies solely on semantic information, resulting in a lower spatial score but an increased emphasis on semantic relevance. Similarly, when the dense semantic module is removed, the Spatial Dense Pass Rate reaches its highest value, as the model depends entirely on spatial modules for retrieval. However, this also leads to a lower Semantic Pass Rate, as semantic information is no longer incorporated into the ranking process. Overall, our spatial-RAG framework integrates both spatial and semantic information, leveraging multi-objective optimization to provide more effective and well-balanced spatial object recommendations.
5.4 Case Study
An example of polyline search is given in Figure 3. Spatial-RAG effectively identifies the user’s intent within a noisy query, detecting that the user needs to travel between two locations (Yankee Stadium and Hilton New York) and is specifically looking for dinner recommendations (”any recos for dinner that evening - something simple”). Based on this, Spatial-RAG classifies the spatial operation as a ”point within polyline” query and applies a 500m buffer distance. Within this region, it filters restaurant spatial objects (shown in Fig. 4). In contrast, traditional methods, such as distance-based approaches, typically generate a buffer zone around a single point, which may not necessarily include POIs along the user’s travel route. Our framework provides a more context-aware understanding of the user’s spatial intent, enabling more precise and relevant recommendations based on both location and user preferences.
Figure 4: An example of a ”point within polyline” query. The user intends to travel between Yankee Stadium and Hilton New York, seeking dinner recommendations along the route. The buffered region (500m) is highlighted, and restaurants are selected.
6 Conclusion
Spatial-RAG enhances LLMs’ spatial reasoning by integrating structured spatial retrieval with natural language understanding, bridging the gap between spatial databases and AI-driven question answering. Our framework improves spatial intelligence, enabling applications in urban planning, tourism, navigation, and geographic QA. Extensive evaluations show that Spatial-RAG outperforms existing methods, highlighting its potential to advance spatial analysis, autonomous navigation, and augmented reality. Beyond improving spatial reasoning, Spatial-RAG lays the foundation for future multimodal models that
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Dense Semantic Retrieval and Ranking
In the previous section, we derived the spatial candidate set $C_{s}$ and the spatial relevance score $f_{s}$. Now, we focus on obtaining the semantic candidate set $C_{k}$ and the semantic relevance score $f_{k}$
Given a query $q$ , we define the semantic candidate set $C_{k}(q)$ as:
$C_{k}(q)=\{y\mid c_{k}(y,q)\leq 0,\forall c_{k}\in\mathcal{C}_{k}(q)\},$ (15)
where:
  • $c_{k}(y,q)$ is a constraint function that filters out spatial objects not satisfying the semantic intent of the query.
  • $\mathcal{C}_{k}(q)$ is the set of all semantic constraints (e.g., topic matching, category relevance).
Each spatial object is associated with textual descriptions, including names, reviews, and additional metadata. However, these descriptions often contain irrelevant or verbose details that may obscure meaningful information. To address this, we use an LLM-based masking function $\mathcal{M}_{k}$ to remove spatially redundant information and retain only semantically relevant content. The resulting texts are then encoded into a dense embedding space by a text encoder $\mathcal{E}$. Specifically, given a spatial object text description $d_{t}$,user query $q$, the filtered text representation is:
$v_{t,k}=\mathcal{E}(\mathcal{M}_{k}(d_{t}))\quad v_{q,k}=\mathcal{E}(\mathcal{M}_{k}(q)).$ (16)
The semantic relevance score is then computed using cosine similarity:
$f_{k}=\frac{v_{q,k}\cdot v_{t,k}}{\parallel v_{q,k}\parallel\parallel v_{t,k}\parallel}.$ (17)
This score quantifies how well the spatial object aligns with the query’s semantic intent, irrespective of spatial factors.
Appendix B Implementation Details
B.1 Semantic Parsing for Spatial Database Query
For the geometry objects referenced in user queries, Spatial-RAG initially interacts with the spatial database to locate and match the described objects, such as specific points (e.g., a restaurant), roads, or defined areas and subsequently retrieves the pertinent geometrical data. In scenarios where the specified geometrical object does not exist pre-mapped in the database, Spatial-RAG is designed to construct a temporary geometric object. This temporary object serves as a stand-in to facilitate spatial queries based on the user’s descriptive input. This approach allows Spatial-RAG to handle dynamic spatial inquiries efficiently, even when direct matches are not immediately found within the existing database entries. By creating temporary geometrical representations, Spatial-RAG ensures that all spatial queries are processed accurately, maintaining the integrity and effectiveness of the system in delivering precise spatial information and responses.
Functionally, the same outcome might be achieved through different means, for example, searching for a restaurant near a street could involve searching within a buffered polyline or creating a polygon enclosing the polyline and searching within it. Such flexibility in the system implies various methods to achieve the same goal. This flexibility, however, poses a challenge if the LLM is tasked with generating a complete query directly, as it might lead to the production of hallucinatory, incorrect, or inexecutable code due to confusion or excessive complexity in interpreting spatial data. By structuring the process such that the LLM first identifies the geometry, then determines the function in a step-by-step manner, we mitigate the risks associated with generating errant queries.
B.2 Semantic Retrieval
While spatial databases address spatial constraints based on the query and spatial database, the actual scenario may be complex, for instance, a hotel may be far from the airport on the map but provide a shuttle, which makes it spatially more convenient than a hotel closer but do not provide a shuttle.
Each spatial object is accompanied by textual descriptions, such as names and reviews. However, the text often contains verbose and irrelevant details that hinder effective decision-making. Moreover, for areas with a high density of POIs that meet spatial requirements, it becomes impractical to input all the text information into an LLM (Large Language Model). To manage the data volume and improve relevance to specific queries, these descriptions are summarized across two perspectives: spatial factors and user preferences. We utilize an LLM to preprocess and summarize spatial objects’ reviews offline, storing the results in the database for future comparison. Similarly, the user preferences in the query are dynamically extracted during the online processing stage. The textual aspects of the user preferences and reviews are compared using cosine similarity scores.
Appendix C Other Case Studies
Fig. 5 presents additional case studies, illustrating two spatial object selection strategies: filtering spatial objects within a buffered radius around the user and retrieving spatial objects contained within a specific polygon region.
Figure 5: (a) An example of a ”point within a $\epsilon$-mile radius” query (b) An example of a ”point within polygon” query
Appendix D Prompt list
D.1 Prompt: Spatial Information Extraction
Extract spatial information from user queries, including object geometry type (Point, Polyline (Route), or Polygon (Region)), region name, distance, and buffer distance.
1Analyze the following user query and extract spatial information: "{user_query}"
2
3Current location context:
4- Number of location points: {location_count}
5- Multiple points detected: {is_multi_point}
6
7First, determine the spatial query type based on these rules:
81. For single location point ({location_count == 1}):
9 - Use Region-based if query explicitly mentions a region
10 - Otherwise, use Point-based
11
122. For exactly two points ({location_count == 2}):
13 - Use Route-based if query suggests path/route between points
14 - Otherwise, fall back to Point/Region based rules
15
163. For multiple points ({location_count > 2}):
17 - Only use Point-based or Region-based
18
19Query types:
201. Point-based:
21 - For "nearby" or "close": 1km in dense areas
22 - For "walking distance": 2km
23 - For "not too far": 3km
24
252. Route-based:
26 - ONLY available with exactly 2 points
27 - For walking routes: 1000m buffer
28 - For general routes: 2000m buffer
29 - For scenic/exploration: 3000m buffer
30 - Consider terms: "route", "path", "between", "from...to", "along"
31
323. Region-based:
33 - ONLY if query explicitly mentions these regions:
34 Community/Sub-region names: {’, ’.join(region_names[’nta_names’])}
35 Borough names: {’, ’.join(region_names[’boro_names’])}
36 - Do NOT infer regions from landmarks
37
38Return in strict JSON format:
39{
40 "query_type": "point" | "route" | "region",
41 "region": "matched region name or null",
42 "distance_km": number or null,
43 "buffer_distance": number or null,
44}
D.2 Prompt: Semantic Intent Extraction
Extract semantic intent from user queries, including spatial and nonspatial preference.
1Analyze the following user query and extract constraints: "{user_query}"
2
3First, determine the main purpose of the query by identifying key terms and context:
4
5Restaurant (R) keywords and contexts:
6- Direct terms: "restaurant", "food", "eat", "dining", "meal", "cuisine"
7- Food types: "Chinese", "Thai", "Mexican", "Italian", "sushi", etc.
8- Meal times: "breakfast", "lunch", "dinner", "brunch"
9- Dining related: "menu", "dishes", "chef", "reservation"
10- Even if staying at a hotel, if asking about food/dining, its Restaurant (R)
11
12Hotel (H) keywords and contexts:
13- Must be explicitly looking for accommodation
14- Direct terms: "hotel", "stay", "accommodation", "room", "book"
15- Price per night (e.g., "\$200/night")
16- Hotel names (e.g., "Hyatt", "Marriott")
17- Mentioning a hotel as location reference is NOT H type
18
19Attraction (A) keywords and contexts:
20- Direct terms: "visit", "see", "tour", "explore"
21- Places: "museum", "park", "gallery", "theater"
22- Activities: "sightseeing", "show", "performance"
23
24Important rules:
251. Focus on what the user is ASKING FOR, not what they mention
262. If user mentions staying at a hotel but asks about restaurants, type is R
273. If query is about food/dining/restaurants, type must be R
284. Location references (e.g., "near Hotel X") dont determine type
29
30For each constraint type, extract complete sentences that describe the requirements:
31
32
331. Spatial constraints: Where they want to go
34 Example: "near Times Square" or "in the Upper West Side area"
35
362. User constraints: What specific requirements or preferences they have
37 Example: "family-friendly restaurant with reasonable prices around \$30 per person"
38
39Please return strict JSON format without any comments:
40{
41 "type": "R/H/A",
42 "spatial_constraints": "complete sentence describing location requirements or null",
43 "user_constraints": "complete sentence describing user preferences and requirements or null"
44}
D.3 Prompt: Result Reranking
Rerank retrieved location results based on user query constraints.
1As a local recommendation expert, please rank the following places based on user query constraints.
2
3User Query Constraints:
4- Spatial Constraints: {query_constraints[’spatial_constraints’]}
5- User Preferences: {query_constraints[’user_constraints’]}
6
7Candidate Places:
8{json.dumps(places, ensure_ascii=False, indent=2)}
9
10Please analyze how well each place matches the user constraints and return a sorted list of places.
11Return format should be a JSON array containing sorted indices.
12Only return the index array, e.g., [2,0,1,3] means the 3rd place is the best match, followed by 1st, 2nd, and 4th places.
13Note: Must return indices for all places, array length should equal input place count ({len(places)}).

### v4

@@
-Spatial-RAG: Spatial Retrieval Augmented Generation for Real-World Spatial Reasoning Questions
+Spatial-RAG: Spatial Retrieval Augmented Generation for Real-World Geospatial Reasoning Questions
 Abstract
-Spatial reasoning remains challenging for Large Language Models (LLMs), which struggle with spatial data retrieval and reasoning.
-We propose Spatial Retrieval-Augmented Generation (Spatial-RAG), a framework that extends RAG to spatial tasks by integrating sparse spatial retrieval (spatial databases) and dense semantic retrieval (LLM-based similarity).
-A multi-objective ranking strategy balances spatial constraints and semantic relevance, while an LLM-guided generator ensures coherent responses.
-Experiments on a real-world tourism dataset show that Spatial-RAG significantly improves spatial question answering, bridging the gap between LLMs and spatial intelligence.
-RAG, spatial, LLM agent, tourism question answering
+Answering real-world geospatial questions—such as finding restaurants along a travel route or amenities near a landmark—requires reasoning over both geographic relationships and semantic user intent.
+However, Existing large language models (LLMs) lack spatial computing capabilities and access to up-to-date, ubiquitous real-world geospatial data, while traditional geospatial systems fall short in interpreting natural language.
+To bridge this gap, we introduce Spatial-RAG, a Retrieval-Augmented Generation (RAG) framework designed for geospatial question answering.
+Spatial-RAG integrates structured spatial databases with LLMs via a hybrid spatial retriever that combines sparse spatial filtering and dense semantic matching.
+It formulates the answering process as a multi-objective optimization over spatial and semantic relevance, identifying Pareto-optimal candidates and dynamically selecting the best response based on user intent.
+Experiments across multiple tourism and map-based QA datasets show that Spatial-RAG significantly improves accuracy, precision, and ranking performance over strong baselines.
 1 Introduction
-Spatial questioning has long been a fundamental domain (Obe & Hsu, 2021), which adeptly handles a variety of spatial questions, from identifying closest neighbors to detecting line-polygon intersections.
-However, traditional spatial questioning systems rely on specialized spatial query languages, which are vastly different from human language, making them inaccessible to layman users.
-More critically, these systems lack the ability to infer complex spatial and semantic relationships from the nuanced, context-rich nature of human text, limiting their applicability to real-world question-answering scenarios.
-Recent advancements in Large Language Models (LLMs) have transformed many fields within machine learning (ML), particularly in understanding and generating human-like text.
-This progress has inspired early efforts to bridge the gap between spatial questioning and natural language by directly eliciting spatial knowledge from LLMs.
-These efforts span a range of applications, including geographic encyclopedic question answering (Mai et al., 2021, 2020; Chen, 2014; Scheider et al., 2021; Contractor et al., 2021a), geolocalization (Zhou et al., 2024; Haas et al., 2024), and automated High-definition map generation (Li et al., 2022; Liu et al., 2023).
-Despite these advancements, recent studies indicate that LLMs exhibit significant limitations in spatial reasoning (Mai et al., 2024; Roberts et al., 2023), and struggle with even basic spatial tasks, such as geoparsing (Mai et al., 2024) and understanding relative spatial relationships (Majic et al., 2024).
-This gap becomes particularly evident when handling real-world spatial reasoning tasks, such as the one illustrated in Figure 1, where the system must recommend a restaurant along a specific route within a predefined area.
-Figure 1: An example of a real-world spatial reasoning question with nearby spatial objects. Areas that satisfy the spatial constraint are highlighted in blue.
-This paper tackles a novel challenge: augmenting LLMs with spatial reasoning capabilities.
-Specifically, we extend Retrieval-Augmented Generation (RAG) into spatial information retrieval and reasoning, bridging the gap between structured spatial databases and unstructured textual reasoning.
-RAG has demonstrated its effectiveness in knowledge-intensive tasks, such as question answering (QA) (Siriwardhana et al., 2023), by retrieving domain-specific documents to enhance LLM responses.
+Figure 1: A real-world spatial reasoning question with nearby spatial objects. Areas that satisfy the spatial constraint are highlighted in purple.
+Spatial reasoning questions are those that require spatial computing to resolve relationships between objects, positions, or movements in space.
+Extensive research has been conducted on abstract spatial reasoning tasks such as mental rotation, block manipulation, and robot navigation, which rely on simplified, small-scale, and often purely geometric representations, typically addressed using techniques from computer vision and robotics.
+In contrast, geospatial reasoning involves interpreting large-scale, real-world geographic data where spatial information is deeply entangled with rich semantics chen2014parameterized ; mai2021geographic ; kefalidis2024question .
+For example, urban routing decisions depend not only on road geometries but also on attributes such as traffic regulations, land use, and temporal constraints.
+Travel plan recommendations should not only consider minimizing the travel distance, but also maximize the quality of the attractions according to their descriptions and reviews xietravelplanner .
+Geospatial reasoning has a longstanding role in AI research, yet classical methods—such as spatial databases and GIS query systems—lack the ability to effectively interpret users’ natural language questions mai2021geographic .
+On the other hand, large language models (LLMs) exhibit strong linguistic competence but struggle with spatial computing and geospatial grounding mai2024opportunities .
+Recent efforts to bridge this gap have focused on prompt engineering manvigeollm ; gurnee2024language , but these approaches heavily rely on LLMs’ internal knowledge, which remains limited in generalization and spatial reasoning capabilities, significantly suffering from geographic bias faisal2023geographic ; manvi2024large ; wu2024torchspatial , and being susceptible to obsolescence as knowledge evolves.
+Some work has explored fine-tuning LLMs on spatial tasks ji2023evaluating ; manvigeollm ; zhang2024bb , but the resulting models are often tailored to narrow applications, constrained datasets, or specific geographic domains.
+Therefore, there remains a critical need for a general-purpose geospatial reasoning framework that synergizes semantic understanding and spatial computation while ensuring access to real-world, vast, fast-changing, and complex geospatial data.
+To fill this gap, this paper aims to augment LLMs with capabilities of spatial reasoning and accessibility to real-world geospatial data.
+For example, as illustrated in Figure 1, answering the question requires LLMs to elicit and formulate the user’s textual request into the problem of “finding points near the polyline" and solve it based on a geospatial map (database) with semantic information (e.g., customer reviews and location profiles).
+Then, it also requires inferring user intent to select the spatially and semantically preferred candidates.
+Thus, the system must seamlessly integrate structured spatial retrieval with unstructured text-based reasoning, ensuring both spatial accuracy and contextual understanding.
+Specifically, we extend Retrieval-Augmented Generation (RAG) into geospatial information retrieval and reasoning, bridging the gap between structured spatial databases and unstructured textual reasoning.
+RAG has demonstrated its effectiveness in knowledge-intensive tasks, such as question answering (QA) siriwardhana2023improving , by retrieving domain-specific documents to enhance LLM responses.
 However, existing RAG systems primarily focus on retrieving and generating textual content and lack the spatial intelligence required for spatial reasoning tasks, especially tasks that involve understanding and computing complex spatial relationships among geometries, including points, polylines, and polygons.
-Spatial questioning is intrinsically challenging, requiring a synergistic combination of text-based inference and spatial computations.
-For example, as illustrated in Figure 1, answering the question requires LLM to elicit and formulate the user’s textual request into the problem of ”finding points near the polyline” and solve it based on a spatial map (database).
-Then, it also requires inferring user preference to select the spatially and semantically preferred candidates.
-Thus, the system must seamlessly integrate structured spatial retrieval with unstructured text-based reasoning, ensuring both spatial accuracy and contextual understanding.
-To achieve this, we introduce Spatial Retrieval-Augmented Generation (Spatial-RAG), a novel framework that unifies text-guided spatial retrieval with spatially aware text generation.
-Specifically, to identify spatially relevant candidate answers, we propose a novel spatial hybrid retrieval module combining sparse and dense retrievers.
+In this paper, we introduce Spatial Retrieval-Augmented Generation (Spatial-RAG), a new framework that unifies text-guided spatial retrieval with spatially aware text generation under multi-objective optimization scenario.
+Specifically, to identify spatially relevant candidate answers, we propose a novel spatial hybrid retrieval module synergizing spatial sparse and dense retrievers.
 To rank the candidates and generate the final answers, we propose to fuel the generator with retrieved results on the Pareto frontier based on a spatial and semantic joint ranking strategy.
 Our contributions are summarized as follows:
-  • A generic Spatial RAG framework: We introduce spatial-RAG, the first framework that extends RAG to spatial question answering, to tackle a broad spectrum of spatial reasoning tasks, such as geographic recommendation, spatially constrained search, and contextual route planning.
+  • A generic spatial RAG framework: We introduce spatial-RAG, the first framework that extends RAG to geospatial question answering, to tackle a broad spectrum of spatial reasoning tasks, such as geographic recommendation, spatially constrained search, and contextual route planning.
 Our approach seamlessly integrates spatial databases, LLMs, and retrieval-based augmentation, enabling effective handling of complex spatial reasoning questions directly within the familiar operational paradigm of LLMs.
-  • Sparse-dense spatial hybrid retriever: We propose a hybrid retrieval mechanism that combines sparse retrieval (SQL-based structured queries) with dense retrieval (LLM-powered semantic matching).
-This dual approach ensures that retrieved results align both spatially and semantically with the user’s query, significantly improving retrieval accuracy in spatial contexts.
-  • Multi-objective guided spatial text generator: To handle both spatial constraints and textual reasoning in the spatial question-answering task, we introduce a multi-objective optimization framework that dynamically balances trade-offs between spatial and semantic relevance.
-This ensures that generated responses are both geometrically accurate and linguistically coherent.
-  • Real-World Evaluation: We evaluated our method on a real-world dataset collected from a Tourism website that has user questions and reviews about different spatial entities.
-Experiments on this dataset reveal the ability to handle real-world spatial reasoning questions.
-Through these innovations, Spatial-RAG significantly enhances the spatial reasoning capabilities of LLMs, bridging the gap between structured spatial databases and natural language question answering.
+  • Sparse-dense spatial hybrid retriever: We propose a hybrid retrieval mechanism that combines spatial sparse retrieval (e.g., SQL-based structured queries) with spatial dense retrieval (e.g., LLM-powered semantic matching).
+This dual approach ensures that retrieved results align both spatially and semantically with the user’s query, synergizing spatial computing and geographical text understanding.
+  • Multi-objective guided spatial and semantic text generator: To handle both spatial constraints and semantic intents in the spatial question-answering task, we introduce a multi-objective optimization framework that dynamically balances trade-offs between spatial and semantic relevancy to the user’s query.
+This ensures that the generated responses are both geospatially accurate and linguistically coherent.
+  • Real-world evaluation: We evaluated our method on multiple real-world datasets consisting of user-generated QA pairs about various spatial entities.
+The experiments demonstrate the model’s ability to handle spatial reasoning questions grounded in real-world scenarios.
+Through these innovations, Spatial-RAG significantly enhances the spatial reasoning capabilities of LLMs, bridging the gap between structured spatial databases and natural language QA.
 2 Related work
 2.1 Retrieval augmented generation
-Retrieval-Augmented Generation (RAG) is a hybrid approach that integrates retrieval systems and generative models to enhance factual accuracy and contextual relevance in natural language generation (Fan et al., 2024).
+Retrieval-Augmented Generation (RAG) is a hybrid approach that integrates retrieval systems and generative models to enhance factual accuracy and contextual relevance in natural language generation fan2024survey .
 Unlike conventional language models that rely solely on parametric memory, RAG dynamically retrieves relevant external knowledge before generating a response.
-One of the foundational works in RAG is Lewis et al.
-(2020), where a retrieval module fetches relevant passages from a large-scale knowledge corpus (e.g., Wikipedia), which are then fused with the question context to generate a more informed response.
+In RAG lewis2020retrieval , a retrieval module fetches relevant passages from a large-scale knowledge corpus (e.g., Wikipedia), which are then fused with the question context to generate a more informed response.
 This technique has proven particularly effective in open-domain question answering (QA), fact verification, and context-aware text generation.
-RAG systems have expanded beyond text and document retrieval to incorporate a wide variety of data types (He et al., 2024) — tables, graphs, charts, and diagrams.
+RAG systems have expanded beyond text and document retrieval to incorporate a wide variety of data types he2024g — tables, graphs, charts, and diagrams.
 While RAG has been widely explored, its application in spatial reasoning question answering remains an unexplored research area.
-Existing studies have primarily focused on knowledge-grounded dialogues (Yu et al., 2024) but often struggle with integrating spatial computation into the question-answering process effectively.
-2.2 Spatial Questions
-Spatial questions in domain-specific applications can generally be categorized into two distinct types: 1) Textual Knowledge-based Spatial Questions These are spatial questions that can be answered by traditional QA methods without the need for spatial computation and reasoning (Liétard et al., 2021).
-For example, the question ”What is the population of Los Angeles city?” falls under this category.
-Despite their spatial context, these questions are essentially text-based and, hence, can be effectively addressed using traditional Retrieval-Augmented Generation (RAG) methods (Christmann & Weikum, 2024).
+Existing studies have primarily focused on knowledge-grounded dialogues yu2024llms but often struggle with integrating spatial computation into the question-answering process effectively.
+2.2 Geospatial question and answering
+Spatial questions in domain-specific applications can generally be categorized into two distinct types: 1) Textual knowledge-based spatial questions These are spatial questions that can be answered by traditional QA methods without the need for spatial computation and reasoning lietard2021language .
+For example, the question "What is the population of Los Angeles city?" falls under this category.
+Despite their spatial context, these questions are essentially text-based and, hence, can be effectively addressed using traditional Retrieval-Augmented Generation (RAG) methods christmann2024rag .
 2) Spatial reasoning questions This category encapsulates spatial questions that demand a model’s capability to comprehend and reason with spatial data and spatial relationships.
-A common example is a model being presented with textual information describing the spatial relationships among multiple objects (Li et al., 2024).
-An example question could be, ”What is the position of object A relative to object B?”, where objects A and B are locations or entities specified on the map.
+A common example is a model being presented with textual information describing the spatial relationships among multiple objects li2024advancing .
+An example question could be, "What is the position of object A relative to object B?", where objects A and B are locations or entities specified on the map.
 Resolving such queries requires a profound understanding of spatial concepts and robust reasoning skills, which largely depend on the model’s training to handle spatial data.
-Several studies (Mai et al., 2024; Roberts et al., 2023) have investigated the capacity of LLMs to understand spatial concepts, yet these models often struggle with accurate reasoning even after fine-tuning.
-Other research (Li et al., 2023) has attempted to enhance this ability by converting geolocation coordinates into addresses to enrich the semantic context.
+Several studies (mai2024opportunities, ; roberts2023gpt4geo, ) have investigated the capacity of LLMs to understand spatial concepts, yet these models often struggle with accurate reasoning even after fine-tuning.
+Other research li2023geolm has attempted to enhance this ability by converting geolocation coordinates into addresses to enrich the semantic context.
 However, these improvements tend to be marginal and are mostly limited to straightforward reasoning tasks like describing positions.
@@
 3 Problem formulation
-In this study, our primary focus is Spatial Reasoning Questions. We formulate the problem as follows: Given a question $q$, the system aims to generate an answer $y$,
+In this study, our primary focus is Geospatial Reasoning Questions.
+We formulate the problem as follows: Given a query $q$, the system aims to generate an answer $y^{*}$, which maximizes the joint spatial and semantic scores while satisfying the spatial constraints in the query $q$.
+For example, in Figure 1, the desired answer is a restaurant that satisfies a spatial constraint—it must be within walking distance of the route.
+Additionally, it should ideally be located along the street (spatial score) and preferably offer waffle fries (semantic score).
+This problem can be formulated as the following multi-objective optimization problem:
 $y^{*}$ $=\arg\max_{y}\ \lambda_{s}^{T}f_{s}(q,y)+\lambda_{k}^{T}f_{k}(q,y)$ (1)
-s.t. $y\in C_{s}(q),\quad y\in C_{k}(q),$
-$\lambda_{s}\geq 0,\quad\lambda_{k}\geq 0,$
-$\mathbf{1}^{T}\lambda_{s}+\mathbf{1}^{T}\lambda_{k}=1,$
+s.t. $y\in C_{s}(q),\quad y\in C_{k}(q),\;\lambda_{s}\geq 0,\lambda_{k}\geq 0,$ $\mathbf{1}^{T}\lambda_{s}+\mathbf{1}^{T}\lambda_{k}=1,$
 where $f_{s}\in\mathbb{R}^{d_{s}}$ is the spatial relevance score vector, $f_{k}\in\mathbb{R}^{d_{k}}$ is the semantic relevance score vector, $C_{s}$ is the spatial candidate set that satisfies the spatial constraints of the question, $C_{k}$ is the semantic candidate set that satisfies the semantic constraints of the question, $\lambda_{s},\lambda_{k}$ are the spatial weights and semantic weights, respectively, $y^{*}$ is the optimal answer, $\mathbf{1}^{T}\lambda_{s}+\mathbf{1}^{T}\lambda_{k}=1$ ensures a normalized trade-off.
-To solve this problem, there are three questions to answer: 1) How to acquire the spatial candidate set $C_{s}$?
-2) How to synergize the spatial and texts by evaluating $f_{s}(q,y)$?
-3) How to trade off the spatial and semantic aspects?
-4 Methodology
+Note that this problem can have multiple valid solutions depending on the trade-off parameters $\lambda_{s}$ and $\lambda_{k}$, forming a Pareto frontier.
+Existing approaches are unable to solve this problem effectively, as it demands a synergistic capability in both geospatial and semantic reasoning.
+Specifically: 1) it requires accurately determining whether a candidate $y$ satisfies the spatial constraints expressed in the query $q$.
+2) it involves ranking candidates based on both spatial and semantic relevance, which are interrelated yet provide complementary signals; and 3) it must ensure that no high-quality answers are overlooked under different trade-offs between spatial and semantic aspects.
+Current methods, which typically excel in either geospatial reasoning (e.g., spatial databases) or semantic reasoning (e.g., large language models and their variants), cannot address all of these requirements without substantial effort to integrate both capabilities seamlessly.
+4 Spatial-RAG for geospatial reasoning questions
 Figure 2: Illustration of the proposed Spatial-RAG framework.
 4.1 Overview
-Our proposed framework, Spatial Retrieval-Augmented Generation (Spatial-RAG), is illustrated in Figure 2.
-Spatial-RAG consists of three key stages: First, to construct the spatial candidate set $C_{s}$, the system must precisely define spatial constraints and then retrieve spatial objects that satisfy them.
-As depicted in Figure 2 (Sparse Spatial Retrieval), we achieve this by parsing the input natural language questions into a spatial SQL query, which will be executed on the spatial database to efficiently retrieve relevant spatial objects from the database.
+Our Spatial-RAG is illustrated in Figure 2, which consists of three key stages: First, to construct the spatial candidate set $C_{s}$, the system must precisely define spatial constraints and then retrieve spatial objects that satisfy them.
+As depicted in Figure 2 (a) Sparse spatial retrieval, we achieve this by parsing the input natural language questions into a spatial SQL query, which will be executed on the spatial database to efficiently retrieve relevant spatial objects from the database.
 This process is detailed in Section 4.2.
 Second, to effectively compute spatial relevance $f_{s}(q,y)$ while integrating textual information, we propose a hybrid spatial retrieval scheme.
-As shown in Figure 2 (Dense Retrieval Module), this method combines sparse spatial relevance scores from the database with dense semantic similarity scores from text embeddings.
+This method combines sparse spatial relevance scores from the database with dense semantic similarity scores from text embeddings.
 This enables the system to rank retrieved spatial objects based on their spatial relevance to the input question, as detailed in Section 4.3.
-Third, given both spatial and semantic constraints, we formulate a multi-objective optimization problem to balance these factors.
-The system computes the Pareto front of candidate answers, and the LLM dynamically trades off among these solutions to generate an optimal response.
+Third, given both spatial and semantic scores, we formulate a multi-objective optimization problem to balance these factors.
+The system computes the Pareto frontier of candidate answers, and the LLM dynamically trades off among these solutions to generate an optimal response.
 This step is covered in Section 4.4.
@@
 The answer to a spatial reasoning question must meet specific spatial constraints.
-The spatial candidate set $C_{s}(q)$ consists of all possible answers $y$ that satisfy a set of spatial constraints $\mathcal{C}_{s}(q)$.
-Formally, we define:
+The spatial candidate set $C_{s}(q)$ consists of all possible answers $y$ that satisfy a set of spatial constraints $\mathcal{C}_{s}(q)$:
 $C_{s}(q)=\{y\mid{c}_{s}(y,q)\leq 0,\forall{c}_{s}\in\mathcal{C}_{s}(q)\},$ (2)
@@
 $\epsilon$ is the set of numerical parameters and spatial relationships governing the spatial constraint (e.g., distance threshold, topological relations).
-Given the diversity and potentially complex nature of these constraints, Large Language Models (LLMs) often struggle to directly construct a complete and executable spatial query from user input.
-To bridge this gap, we structure the spatial query incrementally, allowing the LLM to systematically populate the required components.
+Given the diversity and complex nature of these constraints, LLMs often struggle to directly construct a complete and executable spatial query from user input.
+To bridge this gap, we parse the spatial query incrementally, allowing LLMs to systematically populate the required components.
 Our approach follows three key steps: 1) Geometry recognition: Identify and extract the reference spatial objects $G_{r}$ and candidate target spatial objects $G_{t}$ from the user’s input and extract their spatial footprints – geometries.
@@
   • Point: $\mathcal{G}_{\text{point}}=\{g\mid g\in\mathbb{R}^{2},\dim(g)=0\}$ This category includes single points and multipoints, representing locations with negligible area.
-Examples include stop signs, address points, and a user’s current location.
+Examples include stop signs, points of interest, and a user’s current location.
 In spatial databases, these entities are typically represented as the ’Point’ geometry type.
@@
 The complexity of a spatial query depends on the types of spatial footprints of objects involved.
-For simpler queries, such as ”finding the nearest bus stop from a given location”, only point geometries are required, and the spatial candidate set is
+For simpler queries, such as "finding the nearest bus stop from a given location", only point geometries are required, and the spatial candidate set is
 $C_{s}=\{g|g\in\mathcal{G}_{\text{point}},d(g,g_{\text{point}})<\epsilon\}$ (5)
 where $g_{\text{point}}\subseteq\mathcal{G}_{\text{point}}$ represents a point object (e.g., given location), $\epsilon$ is the distance threshold.
-For more complex queries, such as ”I will walk from home to the university campus along 7th Street and Jones Street; please recommend a café where I can buy breakfast on my walk.”, multiple geometry types must be considered, and the spatial candidate set is
+For more complex queries, such as "I will walk from home to the university campus along 7th Street and Jones Street; please recommend a café where I can buy breakfast on my walk.", multiple geometry types must be considered, and the spatial candidate set is
 $C_{s}=\{g|g\in\mathcal{G}_{\text{point}},g\in B(g_{\text{polyline}},\epsilon)\cup g_{\text{polygon}}\}$ (6)
@@
 By leveraging these mathematical formulations, the system effectively translates spatial reasoning tasks into executable queries, facilitating robust spatial intelligence within the LLM framework.
-4.3 Hybrid Spatial Objects Ranking
-The spatial relevance score $f_{s}$ consists of two components: a score derived from sparse spatial retrieval from the spatial database and a score from spatial dense retrieval based on text similarity between the question and the spatial descriptions of candidate objects.
+4.3 Hybrid spatial relevance scoring
+The spatial relevance score $f_{s}$ consists of two components: a score derived from sparse spatial retrieval from the spatial database and a score from dense spatial retrieval based on text similarity between the question and the spatial descriptions of candidate objects.
 Formally, we define:
-$f_{s}=\lambda_{s}f_{s}^{\text{sparse}}+\lambda_{d}f_{s}^{\text{dense}},$ (8)
-where $\lambda_{s}$ and $\lambda_{d}$ are weighting coefficients controlling the contribution of each score.
+$f_{s}=\lambda_{p}f_{s}^{\text{sparse}}+\lambda_{d}f_{s}^{\text{dense}},$ (8)
+where $\lambda_{p}$ and $\lambda_{d}$ are weighting coefficients controlling the contribution of each score.
 4.3.1 Sparse spatial relevance scoring
@@
 We leverage an LLM to extract key spatial attributes from user queries and compare them with the descriptions of candidate objects.
-Extracting Spatial Requirements
-Given a user query $q$ and a set of text descriptions $d_{t}$ for spatial objects $G_{t}$, we extract the relevant spatial content via an attention-based masking function:
+Extracting spatial requirements and ranking via cosine similarity
+Given a user query $q$ and a set of text descriptions $d_{t}$ for spatial objects $G_{t}$, we extract the spatial requirements via an attention-based masking function:
 $v_{q,s}=\mathcal{E}(\mathcal{M}_{s}(q)),\quad v_{t,s}=\mathcal{E}(\mathcal{M}_{s}(d_{t})),$ (10)
-where $v_{q,s}$ and $v_{t,s}$ are dense vector representations of spatial features, and $\mathcal{M}$ is the extraction function mapping input text to a spatial related text, $\mathcal{E}$ is the text encoder.
-Ranking via Cosine Similarity
-The relevance score is computed via cosine similarity:
-$f_{s}^{\text{dense}}=\frac{v_{q,s}\cdot v_{t,s}}{\parallel v_{q,s}\parallel\parallel v_{t,s}\parallel}.$ (11)
-4.3.3 Hybrid Ranking as a Generalized Model
-We can demonstrate that hybrid ranking generalizes both sparse and dense ranking approaches:
-  • Sparse-Only Case: If $\lambda_{d}=0$ , then $f_{s}=\lambda_{s}f_{s}^{\text{sparse}}$, reducing to a purely distance-based ranking.
-  • Dense-Only Case: If $\lambda_{s}=0$ , then $f_{s}=\lambda_{d}f_{s}^{\text{dense}}$, reducing to a purely semantic-based ranking.
-  • Hybrid Case (General): If both weights are nonzero, hybrid ranking benefits from both explicit spatial constraints and implicit semantic relevance, leading to a more comprehensive ranking mechanism.
-This formulation ensures that hybrid ranking outperforms any single-ranking approach by capturing both spatial proximity and semantic alignment.
+where $v_{q,s}$ and $v_{t,s}$ are dense vector representations of spatial features from query $q$ and text descriptions $d_{t}$.
+$\mathcal{M}_{s}$ is a function mapping input text to a spatial related text.
+$\mathcal{E}$ is the text encoder.
+The dense spatial relevance score $f_{s}^{\text{dense}}$ is computed via cosine similarity between two vectors.
+4.3.3 Hybrid spatial scoring as a generalized model
+We can demonstrate that hybrid spatial scoring generalizes both sparse and dense approaches:
+  • Sparse-only case: If $\lambda_{d}=0$ , then $f_{s}=\lambda_{p}f_{s}^{\text{sparse}}$, reducing to a purely distance-based ranking.
+  • Dense-only case: If $\lambda_{p}=0$ , then $f_{s}=\lambda_{d}f_{s}^{\text{dense}}$, reducing to a purely semantic-based ranking.
+  • Hybrid case (General): If both weights are nonzero, hybrid spatial scoring benefits from both explicit constraints and implicit relevance.
+This formulation ensures that hybrid spatial scoring outperforms any single-scoring approach.
 4.4 Multi-objective generation
-The semantic candidate set $C_{k}$ and the semantic relevance score $f_{k}$ are calculated based on dense vector similarity, we put the details in appendix A.
-After all the scores and candidate sets are acquired, the problem becomes a multi-objective optimization problem since each perspective (spatial and semantic) contributes independently.
-4.4.1 Pareto Front Computation
+The semantic candidate set $C_{k}$ and the semantic relevance score $f_{k}$ are also calculated based on vector similarity, we put the details in Appendix A.
+After all the scores and candidate sets are acquired, the problem becomes a multi-objective optimization problem since each perspective (spatial and semantic) contributes from different aspects.
+4.4.1 Spatial-semantic Pareto frontier computation
 Given the spatial and semantic relevance scores, our goal is to identify the Pareto-optimal candidates that achieve the best trade-off between these objectives.
-A candidate $y$ is Pareto-optimal if no other candidate dominates it in both spatial and semantic relevance.
-Formally, the Pareto front $P(q)$ is defined as:
+A candidate $y$ is spatial-semantic Pareto-optimal if no other candidate dominates it in both spatial and semantic relevance.
+Formally, the Pareto frontier $P(q)$ is defined as:
 […]
-$f_{s}(q,y^{\prime})\geq f_{s}(q,y)\ \text{and}\ f_{k}(q,y^{\prime})\geq f_{k}(q,y),$
-[…]
+$f_{s}(q,y^{\prime})\geq f_{s}(q,y)\ \text{and}\ f_{k}(q,y^{\prime})\geq f_{k}(q,y),\text{with at least one strict inequality}\}.$
 This ensures that each candidate in $P(q)$ is non-dominated, meaning no other candidate is strictly better in both spatial and semantic relevance.
 4.4.2 LLM-based trade-Off decision
-Once the Pareto front $P(q)$ is determined, we use an LLM to dynamically balance the trade-offs between spatial constraints and semantic preferences based on the context of the user query.
-Specifically, the LLM receives the user query, sparse spatial relevance scores, and spatial object descriptions as input:
+Once the Pareto frontier $P(q)$ is determined, we use an LLM to dynamically balance the trade-offs between spatial constraints and semantic preferences based on the context of the user query.
+Specifically, the LLM receives the user query $q$, sparse spatial relevance scores $f_{s}^{\text{sparse}}$, and spatial object descriptions $d_{y}$ as input:
 $I=\left\{q,\left(f_{s}^{\text{sparse}}(q,y)\ ,d_{y}\right)\ ,\forall y\in P(q)\right\}.$ (12)
@@
 By structuring decision-making into discrete steps (candidate filtering $\rightarrow$ Pareto selection $\rightarrow$ trade-off balancing $\rightarrow$ response generation), the LLM avoids generating infeasible or illogical results.
-This structured approach maximizes accuracy and usability, ensuring that the system’s final response aligns closely with the original user intent.
 5 Experiment
 5.1 Experiment setting
-5.1.1 Datasets
-TourismQA (Contractor et al., 2019) dataset contains user questions crawled from TripAdvisor posts from 50 cities around the world.
-Reviews of restaurants, attractions, and hotels for each city are crawled from travel forums and hotel booking websites.
-We select two two popular tourist cities, New York City and Miami, to evaluate the performance of different methods.
-For New York City, the original dataset contains information on 9,470 Points of Interest (POIs), with a total of 17,448 QA pairs.
-For Miami, the dataset includes 2,640 POIs and 133 QA pairs.
-Our preprocessing steps on the dataset include removing POIs with empty review information and eliminating duplicate QA pairs.
-5.1.2 Evaluation Metrics
-We assess the experimental results from multiple dimensions: 1) Delivery Rate: Following (Xie et al., 2024), this metric assesses whether the method can deliver a result successfully.
-2) Spatial Sparse Pass Rate: This metric evaluates whether the parsed spatial query is correct.
-3) Spatial Dense Pass Rate: This metric evaluates whether the answer satisfies the spatial-related semantic constraints in the question.
-4) Semantic Pass Rate: This metric evaluates whether the answer meets the semantic constraints in the question.
-We use GPT-4o to evaluate the last three metrics.
-5.1.3 Models for comparison
-City Model Delivery Rate Spatial Sparse Pass Rate Spatial Dense Pass Rate Semantic Pass Rate
-NYC SD 100 - 57.1 32.1
-TE 99.6 - 48.6 54.6
-ST 100 - 53.6 55.9
-Naive RAG 99.8 - 52.0 54.5
-GeoLLM 99.8 - 69.5 42.8
-Spatial-RAG (GPT-3.5-Turbo) 87.2 67.0 64.4 47.4
-Spatial-RAG (GPT-4-Turbo) 86.1 65.0 71.6 50.1
-Miami SD 100 - 36.8 22.4
-TE 100 - 28.9 39.5
-ST 100 - 26.3 42.1
-Naive RAG 100 - 31.6 40.8
-GeoLLM 100 - 52.6 36.8
-Spatial-RAG (GPT-3.5-Turbo) 86.8 75.8 57.6 43.9
-Spatial-RAG (GPT-4-Turbo) 86.8 81.8 51.5 45.5
-Table 1: Performance comparison of models in New York City (NYC) and Miami.
-The framework is deployed on GPT-4-Turbo and GPT-3.5-Turbo and compared its performance with several baseline models, including SD, TE, ST, Naive RAG, and GeoLLM.
-Model Delivery Rate Spatial Sparse Pass Rate Spatial Dense Pass Rate Semantic Pass Rate
-Spatial-RAG (GPT-4-Turbo) 86.1 65.0 71.6 50.1
-w/o sparse spatial 98.9 - 53.3 78.4
-w/o dense spatial 89.3 61.5 68.4 49.6
-w/o dense semantic 85.9 72.8 75.9 34.8
-Table 2: Different modules of the proposed framework are removed for ablation study.
-We used GPT-3.5-Turbo and GPT-4-Turbo as the LLM in our framework to assess the impact of LLM capability on performance and compared it against the following four methods: Sort-by-distance (SD) (Contractor et al., 2021b): This method ranks the candidate spatial objects based on their distance to the reference objects in the spatial question.
-Text embedding (TE) (Cakaloglu et al., 2020): This greedy method minimizes the distance between the vector embeddings of the text description of the reference object and the target object.
-Spatial-text (ST): This approach computes the embeddings of the user’s question and compares the similarity between the question embedding and the text description embedding of the target object.
+Table 1: Overview of Datasets
+Dataset #POIs #QA Pairs
+TourismQA-NYC 9,470 17,448
+TourismQA-Miami 2,640 133
+MapQA-ADJ 92,415 50
+MapQA-AME 92,415 231
+Datasets.
+We conduct our experiments on four datasets, including TourismQA-NYC Dataset and TourismQA-Miami Dataset datasets, which contains user questions about points of interest (POIs) in real-world cities.
+The questions are crawled from TripAdvisor posts, while the reviews of restaurants, attractions, and hotels are collected from travel forums and hotel booking websites.
+For both datasets, we apply similar preprocessing steps, including removing POIs with missing review information and eliminating duplicate QA pairs.
+MapQA-Adjacent Dataset (MapQA-ADJ) features questions such as “What [amenity] is adjacent to [location]?”, emphasizing topological relationships.
+MapQA-Amenities_Around_Specific Dataset (MapQA-AME) focuses on proximity-based queries like “What are the [amenity] within 50m of [location]?” The number of POIs and QA pairs for each dataset is summarized in Table 1.
+Evaluation metrics.
+To assess the recommendation performance of different methods, we evaluate each model using five widely adopted metrics: Precision@k (The proportion of the top-k returned items that are relevant to the query), Recall@k (The proportion of all relevant items for a query that appear within the top-k results), F1@k (The harmonic mean of Precision@k and Recall@k), and NDCG@k (Normalized Discounted Cumulative Gain up to rank k), where $k\in\{1,3,5,10\}$.
+Table 2: Performance comparison of models in TourismQA-NYC and TourismQA-Miami.
+Dataset Method Precision Recall F1 NDCG
+@1 @3 @5 @10 @1 @3 @5 @10 @1 @3 @5 @10 @1 @3 @5 @10
+TourismQA-Miami GeoLLM 0.3158 0.2719 0.2368 0.2053 0.0365 0.0945 0.1246 0.1824 0.0591 0.1094 0.1311 0.1623 0.3158 0.3200 0.3268 0.3600
+Naive RAG 0.2973 0.2252 0.2054 0.1973 0.0814 0.1124 0.1491 0.2301 0.1005 0.1099 0.1310 0.1642 0.2973 0.2753 0.2849 0.3403
+TE 0.3421 0.3070 0.2316 0.1842 0.0567 0.1544 0.1821 0.2231 0.0781 0.1558 0.1515 0.1543 0.3421 0.3766 0.3890 0.4748
+SD 0.1053 0.0351 0.0211 0.0105 0.0075 0.0075 0.0075 0.0075 0.0138 0.0119 0.0105 0.0082 0.1053 0.1053 0.1053 0.1053
+ST 0.3947 0.1316 0.0789 0.0395 0.0873 0.0873 0.0873 0.0873 0.1115 0.0730 0.0569 0.0386 0.3947 0.3947 0.3947 0.3947
+Spatial-RAG(GPT3.5-Turbo) 0.5455 0.4141 0.4000 0.3152 0.1081 0.2072 0.2972 0.3732 0.1540 0.2262 0.2765 0.2817 0.5455 0.492 0.5276 0.5440
+Spatial-RAG (GPT4-Turbo) 0.5152 0.4242 0.3758 0.2939 0.1026 0.2037 0.2841 0.3565 0.1454 0.2246 0.2619 0.2656 0.5152 0.4852 0.4983 0.5079
+TourismQA-NYC GeoLLM 0.3650 0.3292 0.3020 0.2725 0.0168 0.0433 0.0614 0.1033 0.0311 0.0688 0.0907 0.1328 0.3650 0.3626 0.3708 0.4248
+Naive RAG 0.4923 0.4447 0.4245 0.4181 0.0214 0.0556 0.0889 0.1678 0.0399 0.0932 0.1345 0.2167 0.4923 0.4551 0.4528 0.5058
+TE 0.2250 0.2433 0.2395 0.2325 0.0101 0.0330 0.0506 0.0934 0.0189 0.0541 0.0767 0.1199 0.2250 0.2873 0.3533 0.5142
+SD 0.2425 0.2492 0.2455 0.2402 0.0103 0.0314 0.0520 0.0945 0.0193 0.0520 0.0758 0.1194 0.2425 0.2694 0.2926 0.3624
+ST 0.4725 0.4608 0.4460 0.4323 0.0238 0.0644 0.1026 0.1846 0.0433 0.1026 0.1475 0.2281 0.4725 0.4686 0.4745 0.5296
+Spatial-RAG (GPT3.5-Turbo) 0.4611 0.4352 0.4144 0.4098 0.0188 0.0485 0.0751 0.1493 0.0350 0.0828 0.1188 0.2000 0.4611 0.4545 0.4534 0.5097
+Spatial-RAG (GPT4-Turbo) 0.5665 0.4875 0.4555 0.4251 0.0274 0.0611 0.0908 0.1691 0.0483 0.0996 0.1384 0.2153 0.5665 0.5174 0.5065 0.5574
+Table 3: Performance comparison of models on MapQA-ADJ and MapQA-AME.
+Dataset Metric Methods
+SD TE ST Naive RAG GeoLLM Spatial-RAG (GPT-3.5-Turbo) Spatial-RAG (GPT-4.0-Turbo)
+MapQA-ADJ Precision 0.4600 0.4200 0.5600 0.3455 0.4558 0.5467 0.5057
+Recall 0.4200 0.3750 0.5150 0.6170 0.6685 0.5350 0.7819
+F1 0.4333 0.3880 0.5280 0.4067 0.5073 0.5247 0.5625
+NDCG 0.4600 0.4200 0.5600 0.5866 0.6132 0.5800 0.7391
+MapQA-AME Precision 0.4323 0.2751 0.5714 0.2452 0.3729 0.5566 0.6645
+Recall 0.3923 0.2351 0.5173 0.6075 0.8333 0.7446 0.9072
+F1 0.4054 0.2482 0.5346 0.3263 0.4744 0.6100 0.7298
+NDCG 0.4323 0.2751 0.5714 0.6018 0.8199 0.7264 0.8719
+Table 4: Ablation results.
+Method Precision Recall F1 NDCG
+@1 @3 @5 @10 @1 @3 @5 @10 @1 @3 @5 @10 @1 @3 @5 @10
+Spatial-RAG(GPT4-Turbo) 0.5665 0.4875 0.4555 0.4251 0.0274 0.0611 0.0908 0.1691 0.0483 0.0996 0.1384 0.2153 0.5665 0.5174 0.5065 0.5574
+w/o sparse spatial 0.3807 0.3545 0.3503 0.3307 0.0147 0.0431 0.0701 0.1305 0.0276 0.0701 0.1058 0.1667 0.3807 0.3608 0.3570 0.3455
+w/o dense spatial 0.5569 0.4954 0.4628 0.4295 0.0263 0.0606 0.0919 0.1679 0.0466 0.0986 0.1392 0.2132 0.5569 0.5240 0.5132 0.5587
+w/o dense semantic 0.4986 0.4311 0.4165 0.3806 0.0240 0.0553 0.0844 0.1499 0.0421 0.0883 0.1262 0.1896 0.4986 0.4679 0.4739 0.5219
+Scratch 0.5392 0.4935 0.4706 0.4343 0.0236 0.0535 0.0829 0.1574 0.0435 0.0916 0.1323 0.2105 0.5392 0.5105 0.5042 0.5374
+w/o RAG 0.3584 0.3333 0.3028 0.2644 0.0121 0.0323 0.0484 0.0846 0.0231 0.0568 0.0796 0.1191 0.3584 0.3570 0.3653 0.4186
+Models for comparison.
+We used GPT-3.5-Turbo and GPT-4-Turbo as the LLM in our framework and compared them against the following methods: GeoLLM manvigeollm encodes the spatial objects to address and enrich their context by adding spatial information of nearby spatial objects.
+Naive RAG lewis2020retrieval saves all spatial objects’ descriptions in a vector database and retrieves the most relevant objects based on vector similarity.
+Text embedding (TE) cakaloglu2020text is a greedy method minimizing the distance between the vector embeddings of the text description of the reference object and the target object.
+Sort-by-distance (SD) contractor2021joint ranks the candidate spatial objects based on their distance to the reference objects in the spatial question.
+Spatial-text (ST) computes the embeddings of the user’s question and compares the similarity between the question embedding and the text description embedding of the target object.
 Additionally, the object’s location is encoded as a distance score.
-The answer is then determined based on the weighted average of these scores.
-Naive RAG (Lewis et al., 2020): This method saves all spatial objects’ descriptions in a vector database and retrieves the most relevant objects based on vector similarity.
-GeoLLM (Manvi et al., 2024): This method encodes the spatial objects to address and enrich its context by adding spatial information of nearby spatial objects.
-5.2 Main Results
-The main results are summarized in 1 For the NYC dataset, Spatial-RAG demonstrates minimal differences in Delivery Rate when deployed with different LLMs.
-Approximately 86.1% of the questions are successfully processed.
-The failure cases where the framework is unable to generate valid recommendations can be classified into two main categories: Around 12.9% of the questions fail to retrieve any spatial objects from the spatial database, either due to polygon recognition errors or because the specified region in the SQL query lacks relevant spatial objects.
-The remaining 0.9% of cases stem from the LLM’s inability to correctly re-rank the retrieved results during the reranking process.
-Additionally, the performance gap of our framework based on two LLMs in generating spatial SQL queries is insignificant.
-However, GPT-4-Turbo outperforms GPT-3.5-Turbo by approximately 7% in Spatial Dense Pass Rate and also demonstrates a slight advantage in Semantic Pass Rate, and both models perform slightly below the optimal values (by 5% and 8%, respectively).
-This is due to the trade-off made when jointly considering both spatial proximity and the user’s personal preferences.
-Since the SD method returns the nearest spatial object, it achieves a higher spatial dense retrieval score than the TE, ST, and Naive RAG methods.
-However, its exclusive reliance on distance results in inferior performance on other metrics.
-For GeoLLM, as it considers only the names of spatial objects and their distances, it also achieves competitive performance in Spatial Dense Pass Rate.
-In contrast, both the TE and ST methods consider semantic contexts, leading to superior performance in this aspect compared to other models.
-The Naive RAG and ST methods jointly optimize for both spatial dense retrieval and user semantic dense retrieval, resulting in comparable performance between the two models.
-For Miami, which differs significantly from NYC in terms of area and geography, Spatial-RAG also demonstrates strong performance.
-The performance of the baseline models almost remains consistent with the patterns observed in the previous analysis.
-Figure 3: An example of how Spatial-RAG operates: Given a question, 1) Sparse Spatial Retrieval: an LLM parses the natural language question into a spatial SQL query for spatial database, retrieving spatial objects that meet the spatial constraints and the sparse spatial relevance scores.
-2) Question Decomposition and Dense Retrieval: In parallel, Spatial-RAG decomposes the question into spatial and semantic components and compares them with the descriptions of spatial objects to perform dense retrieval, filtering out irrelevant ones.
-3) LLM Reranking: A language agent balances the spatial and semantic aspects to rerank the candidate and generate the final answer.
+The answer is determined based on the average of these scores.
+5.2 Spatial-RAG vs. Baselines
+The comparative performance results across methods and datasets are summarized in Tables 2 and 3.
+Our Spatial-RAG (GPT4-Turbo) consistently surpasses baselines, including Naive RAG, GeoLLM, and template-based methods (ST, TE, SD).
+Specifically, a 19.9% improvement in Precision@1 was achieved by our Spatial-RAG (GPT-4-Turbo) compared to the best baseline model (ST) on TourismQA-NYC.
+It improved NDCG by 32.0% on MapQA-ADJ and by 52.6% on MapQA-AME compared to ST.
+On the MapQA-AME dataset, it notably increased Recall by 75.4% over ST.
+Compared to GeoLLM, Spatial-RAG (GPT-4-Turbo) achieved 95.4% higher Recall@10 on TourismQA-Miami.
+This advantage arises primarily from Spatial-RAG’s structured spatial retrieval pipeline, which leverages a spatial database to accurately retrieve relevant candidates that satisfy the spatial constraints, a capability lacking in embedding-based methods like Naive RAG and purely textual or parametric models (GeoLLM, SD, TE).
+Meanwhile, although ST occasionally leads in single-relationship precision (e.g., Precision@1 on MapQA-ADJ), Spatial-RAG’s GPT4-powered re-ranking provides superior overall performance through nuanced multi-constraint spatial reasoning.
+Among evaluation metrics, Spatial-RAG sees greater gains in Recall and NDCG than in Precision, highlighting that structured spatial queries significantly enhance completeness of retrieved results and the accuracy of ranking order.
+Finally, Spatial-RAG (GPT4-Turbo) consistently outperforms its GPT3.5 counterpart, underscoring GPT4’s advanced reasoning capabilities and improved candidate-ranking accuracy.
+Replacing GPT3.5-Turbo with GPT4-Turbo within the identical retrieval pipeline contributes a further +2 pp on average, indicating that model capacity and spatially constrained retrieval are additive rather than redundant.
+This improvement is especially evident on structured spatial QA tasks such as those in MapQA, which feature well-defined relations like “adjacent” and “within 50m.” In contrast, the performance gap is narrower on noisier, open-domain datasets like TourismQA, suggesting that GPT4’s capacity brings greater benefit when precise spatial reasoning is required.
+The above results establish Spatial-RAG as the new state of the art, delivering statistically robust, across-the-board improvements relative to all prior baselines.
 5.3 Ablation study
-We conducted three ablation studies to assess the importance of different modules in our framework.
-We remove the sparse spatial module, dense spatial module, and dense semantic module, respectively, and denote the ablation versions as w/o sparse spatial, w/o dense spatial, and w/o dense semantic.
-The results are summarized in Table 2.
-It’s found that there is a significant increase in Delivery Rate after removing the sparse spatial module, as this eliminates the need to determine the geometry type.
-In this case, the model relies solely on semantic information, resulting in a lower spatial score but an increased emphasis on semantic relevance.
-Similarly, when the dense semantic module is removed, the Spatial Dense Pass Rate reaches its highest value, as the model depends entirely on spatial modules for retrieval.
-However, this also leads to a lower Semantic Pass Rate, as semantic information is no longer incorporated into the ranking process.
-Overall, our spatial-RAG framework integrates both spatial and semantic information, leveraging multi-objective optimization to provide more effective and well-balanced spatial object recommendations.
-5.4 Case Study
-An example of polyline search is given in Figure 3.
-Spatial-RAG effectively identifies the user’s intent within a noisy query, detecting that the user needs to travel between two locations (Yankee Stadium and Hilton New York) and is specifically looking for dinner recommendations (”any recos for dinner that evening - something simple”).
-Based on this, Spatial-RAG classifies the spatial operation as a ”point within polyline” query and applies a 500m buffer distance.
-Within this region, it filters restaurant spatial objects (shown in Fig.
-4).
-In contrast, traditional methods, such as distance-based approaches, typically generate a buffer zone around a single point, which may not necessarily include POIs along the user’s travel route.
+To validate our technical contributions, we conducted five ablation studies.
+First, we examined the impact of removing individual modules: the sparse spatial module (w/o sparse spatial), dense spatial module (w/o dense spatial), and dense semantic module (w/o dense semantic).
+Next, we evaluated two system-level variations: generating SQL from scratch (Scratch) and removing the RAG component (w/o RAG).
+Table 4 summarizes the ablation results.
+Removing any component leads to a performance decline, highlighting the importance of each module.
+Notably, excluding the sparse spatial component results in the most significant drop, underscoring the value of integrating the spatial database with the LLM.
+The Scratch setting shows only a slight decrease, suggesting that GPT4-Turbo can internally formulate SQL queries when given essential inputs, even without templates.
+The w/o RAG variant performs worst across all metrics, indicating that retrieval-augmented generation is fundamental to the system’s overall effectiveness.
+5.4 Case studies
+Figure 3: Three POI queries with different reference objects.
+The visualizations of three POI queries with different reference objects are given in Figure 3.
+Spatial-RAG effectively identifies the user’s intent within noisy queries, detecting that the user needs to travel near a location, along a route, or within a region.
+Based on this, Spatial-RAG applies proper spatial constraints to filter spatial objects.
+In contrast, traditional methods, such as distance-based approaches, always generate a buffer zone around a single location, which may not necessarily include POIs along the user’s travel route.
 Our framework provides a more context-aware understanding of the user’s spatial intent, enabling more precise and relevant recommendations based on both location and user preferences.
-Figure 4: An example of a ”point within polyline” query.
-The user intends to travel between Yankee Stadium and Hilton New York, seeking dinner recommendations along the route.
-The buffered region (500m) is highlighted, and restaurants are selected.
 6 Conclusion
-Spatial-RAG enhances LLMs’ spatial reasoning by integrating structured spatial retrieval with natural language understanding, bridging the gap between spatial databases and AI-driven question answering.
-Our framework improves spatial intelligence, enabling applications in urban planning, tourism, navigation, and geographic QA.
-Extensive evaluations show that Spatial-RAG outperforms existing methods, highlighting its potential to advance spatial analysis, autonomous navigation, and augmented reality.
-Beyond improving spatial reasoning, Spatial-RAG lays the foundation for future multimodal models that
-Impact Statement
-This paper presents work whose goal is to advance the field of Machine Learning.
-There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
+Spatial-RAG enhances LLMs’ spatial reasoning by integrating structured spatial retrieval with natural language understanding, bridging the gap between spatial databases and LLM-driven question answering.
+Extensive evaluations show that Spatial-RAG outperforms existing methods, highlighting its potential to advance spatial analysis, tourism recommendation, and geographic QA.
+Acknowledgments and Disclosure of Funding
+Use unnumbered first level headings for the acknowledgments.
+All acknowledgments go at the end of the paper before the list of references.
+Moreover, you are required to declare funding (financial activities supporting the submitted work) and competing interests (related financial activities outside the submitted work).
+More information about this disclosure can be found at: https://neurips.cc/Conferences/2025/PaperInformation/FundingDisclosure.
+Do not include this section in the anonymized submission, only in the final paper.
+You can use the ack environment provided in the style file to automatically hide this section in the anonymized submission.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: We discuss the limitations of the work in appendix E
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [N/A]
+  14. Justification: This paper does not involves theoretical result.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: We propose a framework and describe the architecture clearly and fully.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: We will release of code and data once get accepted.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: We specify all the training and test details in the main text and appendix.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [No]
+  34. Justification: We follow the convention in prior works.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: We provided sufficient information on the computer resources in the main text and appendix.
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  40. 9.
+    Code of ethics
+  41. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  42. Answer: [Yes]
+  43. Justification: the research conducted in the paper conformed, in every respect, with the NeurIPS Code of Ethics.
+  44. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  45. 10.
+    Broader impacts
+  46. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  47. Answer: [N/A]
+  48. Justification: there is no societal impact of the work performed
+  49. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  50. 11.
+    Safeguards
+  51.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  52. Answer: [N/A]
+  53. Justification: the paper poses no such risks.
+  54. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  55. 12.
+    Licenses for existing assets
+  56.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  57. Answer: [Yes]
+  58. Justification: We properly credited the creators or original owners of assets (e.g., code, data, models), used in the paper and conformed the license and terms.
+  59. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  60. 13.
+    New assets
+  61. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  62. Answer: [Yes]
+  63. Justification: We communicated the details of the dataset/code/model as part of their submission.
+  64. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  65. 14.
+    Crowdsourcing and research with human subjects
+  66.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  67. Answer: [N/A]
+  68. Justification: does not involve crowdsourcing nor research with human subjects
+  69. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  70. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  71.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  72. Answer: [N/A]
+  73. Justification: Our paper does not involve study participants.
+  74. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  75. 16.
+    Declaration of LLM usage
+  76.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  77. Answer: [No]
+  78. Justification: core method development in this research does not involve LLMs as any important, original, or non-standard components.
+  79. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
 Appendix A Dense Semantic Retrieval and Ranking
@@
 By structuring the process such that the LLM first identifies the geometry, then determines the function in a step-by-step manner, we mitigate the risks associated with generating errant queries.
-B.2 Semantic Retrieval
+B.2 Dense Retrieval
 While spatial databases address spatial constraints based on the query and spatial database, the actual scenario may be complex, for instance, a hotel may be far from the airport on the map but provide a shuttle, which makes it spatially more convenient than a hotel closer but do not provide a shuttle.
@@
 Moreover, for areas with a high density of POIs that meet spatial requirements, it becomes impractical to input all the text information into an LLM (Large Language Model).
-To manage the data volume and improve relevance to specific queries, these descriptions are summarized across two perspectives: spatial factors and user preferences.
+To manage the data volume and improve relevance to specific queries, these descriptions are summarized across two perspectives: spatial requirements and semantic requirements.
 We utilize an LLM to preprocess and summarize spatial objects’ reviews offline, storing the results in the database for future comparison.
-Similarly, the user preferences in the query are dynamically extracted during the online processing stage.
-The textual aspects of the user preferences and reviews are compared using cosine similarity scores.
-Appendix C Other Case Studies
-Fig.
-5 presents additional case studies, illustrating two spatial object selection strategies: filtering spatial objects within a buffered radius around the user and retrieving spatial objects contained within a specific polygon region.
-Figure 5: (a) An example of a ”point within a $\epsilon$-mile radius” query (b) An example of a ”point within polygon” query
-Appendix D Prompt list
-D.1 Prompt: Spatial Information Extraction
+Similarly, the user queries are dynamically extracted during the online processing stage.
+Appendix C Prompt list
+C.1 Prompt: Spatial Information Extraction
 Extract spatial information from user queries, including object geometry type (Point, Polyline (Route), or Polygon (Region)), region name, distance, and buffer distance.
@@
 44}
-D.2 Prompt: Semantic Intent Extraction
+C.2 Prompt: Semantic Intent Extraction
 Extract semantic intent from user queries, including spatial and nonspatial preference.
@@
 44}
-D.3 Prompt: Result Reranking
+C.3 Prompt: Result Reranking
 Rerank retrieved location results based on user query constraints.
@@
 13Note: Must return indices for all places, array length should equal input place count ({len(places)}).
+Appendix D Limitations
+We employ the ChatGPT API for direct LLM-based detection of target and reference objects.
+While this approach provides immediate functionality, future work will investigate fine-tuning a specialized language model to enhance performance for this specific task.
+Appendix E Additional case studies
+E.1 Case 1
+User Query:
+Going to be in Manhattan for a Broadway show this weekend and am looking for a recommendation for a good restaurant in the theatre district that is reasonably priced (\$30 to \$50 per person), good food and not a tourist trap, would love to avoid the hotel restaurants and experience a real NY experience.
+Does such a place exist?
+Someone suggested Carmine’s but they are totally booked.
+(like all kinds of food) Any suggestions would be greatly appreciated - picking one out of the 5000+ listed is a daunting task.
+Model Processing Summary:
+The system first determines that the reference object type is polygon, and correctly identifies the region as Manhattan.
+It detects that the target object type is a restaurant point.
+Based on these constraints, a SQL query is constructed to retrieve 532 candidate locations from the database.
+The spatial requirement is:
+"am looking for a recommendation for a good restaurant in the theatre district"
+The semantic requirements are:
+"reasonably priced (\$30 to \$50 per person), good food and not a tourist trap, avoid hotel restaurants, real NY experience"
+The system computes vector similarities between the query and the reviews of all 532 candidates, considering both spatial and semantic relevance.
+The top 20 results based on combined score are passed to an LLM reranker, and the final top-1 recommendation is:
+West Bank Cafe, 407 W 42nd St, New York City, NY 10036-6808
+Reasoning:
+  • Spatial match: West Bank Cafe is located in the Theatre District of Manhattan, satisfying the user’s regional constraint. Its location aligns well with proximity to Broadway shows.
+  • Semantic alignment: The restaurant is described as having reasonable pricing and good food, with a local, non-touristy atmosphere.
+These attributes align with the user’s preferences for affordability and an authentic New York experience.
+E.2 Case 2
+User Query:
+Going to NY in May, it is my gf’s birthday on our last night – and I would be looking for a nice restaurant to spoil her :)!
+We are staying in Times Square – so would like somewhere close by!
+I would like either an American or Italian cuisine!
+Any advice would be much appreciated.
+Model Processing Summary:
+The model identifies the query as a point spatial search with a radius of 1.0 km around Times Square.
+The target object type is recognized as a restaurant point.
+Based on these criteria, a SQL query returns 119 candidate locations from the database.
+The spatial requirement is:
+"We are staying in Times Square – so would like somewhere close by!"
+The semantic requirements are:
+"it is my gf’s birthday on our last night",
+"I would like either an American or Italian cuisine!"
+The system computes vector similarities based on spatial and semantic constraints. The top 20 candidates are re-ranked by an LLM reranker, and the final top-1 recommendation is:
+Pasta Lovers, 142 W 49th St, New York City, NY 10019-6802
+Reasoning:
+  • Spatial match: Pasta Lovers is located within 1 km of Times Square, satisfying the proximity requirement effectively.
+  • Semantic alignment: It offers Italian cuisine, aligning with the user’s preference for either American or Italian food.
+The atmosphere and reviews suggest a cozy dining experience appropriate for a date, though it is not explicitly described as suitable for a birthday celebration.
+Figure 4: (a) Case 1; (b) Case 2
+E.3 Case 3
+User Query:
+Visiting NYC for the Marathon this weekend.
+Staying in apartment in Hell’s Kitchen area (around 51st and 12th).
+Anyone have good recommendations for a great breakfast restaurant within an 8–10 block radius?
+Not worried about eating healthy the day before—looking for filling, traditional American-style breakfast!
+Also looking for ideas for a good filling dinner for Sunday evening after the race.
+We’re not looking for fine dining, just something good in the area for myself, wife and 16 yo daughter.
+Thanks in advance for any good recommendations.
+Model Processing Summary:
+The model identifies the query as a point spatial query with a radius of 1.5 km, derived from the “8–10 block” constraint.
+The target object type is restaurant point, and the approximate location is Hell’s Kitchen near 51st St and 12th Ave.
+A SQL query retrieves 140 candidates from the database based on this spatial condition.
+The spatial requirement is:
+"Staying in apartment in Hell’s Kitchen area (around 51st and 12th), breakfast restaurant within an 8–10 block radius"
+The semantic requirements include:
+"traditional American-style breakfast", "filling dinner after the race (Sunday evening), just something good in the area for myself, wife and 16 yo daughter."
+The system calculates vector similarity scores based on spatial and semantic requirements, and selects the top 20 results.
+These are re-ranked by an LLM reranker, and the final top-1 recommendation is:
+Galaxy Diner, 665 9th Ave, New York City, NY 10036-3623
+Reasoning:
+  • Spatial match: Galaxy Diner is located within a reasonable walking distance of the user’s location, satisfying the “8–10 block radius” constraint in Hell’s Kitchen.
+  • Semantic alignment: The diner offers traditional American breakfasts like pancakes, eggs, and bacon, aligning with the user’s request for filling, non-healthy food before the race.
+It also accommodates a casual family atmosphere.
+E.4 Case 4
+User Query:
+We are staying in Midtown, so figured it would be easier to go somewhere close by, but I’m not opposed to somewhere a little further away.
+Looking for a restaurant where we can go to dinner and dress up but not have it be a crazy price.
+I understand that a "dress up" place isn’t usually cheap, but something on the lower end of the "dressy" price scale would be great :) Thanks!
+Model Processing Summary:
+The model correctly identifies the reference object type as polygon (Midtown). A SQL query retrieves 97 candidates from the database based on this spatial condition.
+The spatial requirement is:
+"Staying in Midtown, somewhere close by, but not opposed to somewhere a little further away."
+The semantic requirement is:
+"restaurant where we can go to dinner and dress up but not have it be a crazy price, lower end of the ’dressy’ price scale."
+The model calculates spatial and user constraint similarities, ranks the top 20 results, and applies an LLM reranker. The final top-1 recommendation is:
+Nocello, 257 W 55th St, New York City, NY 10019-5232
+Reasoning:
+  • Spatial match: Nocello is within a reasonable walking distance of Midtown, fulfilling the requirement of being nearby but not necessarily adjacent.
+It’s described as a short walk from Restaurant Row and near Broadway, aligning well with the user’s open spatial boundary.
+  • Semantic alignment: The restaurant is described as offering excellent food at a reasonable price, making it suitable for a "dress-up" dinner without the high-end cost.
+This aligns with the user’s goal of finding something elegant but affordable.
+Figure 5: (a) Case 3; (b) Case 4
+E.5 Case 5
+User Query:
+Where are some good places to eat breakfast?
+We are staying in the southern tip of Manhattan near Battery Park, but it doesn’t have to be contained to just that area.
+We are very open to ideas, but are going to be avoiding McDonalds, BK, etc.
+Could be larger restaurants but also would like to visit a few small places and would like to be able to sit and eat outside.
+Model Processing Summary:
+The model correctly identifies the reference object type as polygon (battery park city-lower manhattan). A SQL query retrieves 14 candidates from the database based on this spatial condition.
+The spatial requirement is:
+"staying in the southern tip of Manhattan near Battery Park, but it doesn’t have to be contained to just that area."
+The semantic requirements are:
+"avoiding McDonalds, BK, etc. Could be larger restaurants but also would like to visit a few small places and would like to be able to sit and eat outside."
+The model ranks the top 10 by spatial and user relevance, and performs LLM reranking. The final top-1 recommendation is:
+Stone Street Tavern, 52 Stone St, New York City, NY 10004-2604
+Reasoning:
+  • Spatial match: The restaurant is located a short walk from Battery Park in the financial district, consistent with the user’s desire to explore areas nearby but not strictly limited to Battery Park.
+The cobblestone street setting and access to outdoor space align well with the user’s spatial intent.
+  • Semantic alignment: Stone Street Tavern offers outdoor seating and avoids fast-food chains.
+It provides a casual and local dining experience with bench-style outdoor tables, fitting the user’s interest in both large and small sit-down places for breakfast.
+E.6 Case 6
+User Query:
+We have tickets for the Saturday performance of War Horse at the Vivian Beaumont Theater at Lincoln Center. I would appreciate a recommendation for a reasonable pre-theater dinner.
+Model Processing Summary:
+The model correctly identifies the reference object type as point query, as the user provides a specific point of interest (Vivian Beaumont Theater) without mentioning a formal region.
+A walking distance of 2.0 km is assumed for pre-theater dining.
+A SQL query retrieves 255 candidates from the database based on this spatial condition.
+The spatial requirement is:
+"Vivian Beaumont Theater at Lincoln Center" – implying a location within walking distance
+The semantic requirements are:
+"reasonable pre-theater dinner" — suggesting affordability and suitable timing for a theater schedule.
+The model reranks the top 20 via LLM reranking. The final top-1 recommendation is:
+Bar Boulud, 1900 Broadway, New York City, NY 10023-7004
+Reasoning:
+  • Spatial match: Bar Boulud is located directly across from Lincoln Center, fulfilling the proximity constraint perfectly for a pre-theater dinner.
+  • Semantic alignment: The restaurant is known for accommodating theatergoers and offers timing suitable for pre-show dining.
+However, user reviews mention that while the food and service are excellent, prices may be higher than what the user considers “reasonable”.
+Figure 6: (a) Case 5; (b) Case 6
+E.7 Case 7
+User Query:
+I want to walk the Highline on my forthcoming trip to NY, and wanted some recommendations for a good spot for a sit down lunch somewhere close to one of the exits off the Highline.
+Will be starting north to south.
+Don’t mind what type of food, just nice atmosphere required.
+Have looked at MenuPages but so many restaurants not sure if anyone has any particular favourites?
+Model Processing Summary:
+The model correctly identifies the reference object type as polyline.
+This is inferred from the user’s mention of walking from the north to south along the Highline, with interest in locations near the path (Highline exits).
+Two coordinates representing the start and end of the route are provided.
+The buffer distance is set to 500 meters.
+The spatial requirement is:
+"somewhere close to one of the exits off the Highline."
+The semantic requirement is:
+"sit down lunch, nice atmosphere required."
+The system extracts 52 candidate restaurants along the route, ranks them by spatial and user constraints, and performs LLM reranking on the top 20. The final top-1 recommendation is:
+Barbuto, 775 Washington St, New York City, NY 10014-1748
+Reasoning:
+  • Spatial match: Barbuto is located directly off the Highline and near one of its southern exits, which aligns precisely with the user’s request for proximity to the walking path.
+  • Semantic alignment: The restaurant offers a sit-down experience with a "fancy but cozy" atmosphere, partially fulfilling the user’s request for a "nice atmosphere." However, some reviews describe the food as average, which may not fully satisfy quality expectations.
+E.8 Case 8
+User Query:
+We arrive Sat night at 10:30 (unfortunately AA changed our flight from an earlier one).
+We are spending 1 night at RCA at 142 W 49th between 6th and 7th.
+Sunday we move to Hilton New York at 1335 Ave Of Americas between 53 and 54.
+We have tickets for a Yankee game at 1:05.
+(We will also be going to Dizzy’s Coca Cola at 9PM.) Could you guys give me some tips on the best way to handle the hectic morning?
+Leave our bags at RCA?
+Move them to Hilton for storage so we can get subway to Yankee Stadium?
+Can you give me some recommendations for dinner that evening - something simple.
+Thanks!
+Model Processing Summary:
+The model correctly identifies the reference object type as polyline, given that the user describes movement between two locations: Hilton New York and Yankee Stadium.
+The user’s intention to find a dinner spot along that route aligns with this classification.
+The buffer distance is set to 500 meters.
+"The spatial requirement is:
+Implied need for proximity to Hilton New York or Yankee Stadium for convenience."
+The semantic requirements are:
+"We will also be going to Dizzy’s Coca Cola at 9PM." and "recommendations for dinner that evening - something simple."
+The system identifies 50 restaurant candidates within the route buffer, computes user and temporal similarity, and applies LLM reranking on the top 20. The final top-1 recommendation is:
+Good Enough to Eat, 520 Columbus Ave, Frnt A, New York City, NY 10024-3404
+Reasoning:
+  • Spatial match: This place is within the buffer range, and the restaurant is reasonably accessible and could fit into the evening schedule before Dizzy’s Coca Cola.
+  • Semantic alignment: The user requested a simple dinner.
+The reviews highlight casual meals like pancakes, BLT omelets, and quick service, matching the preference for something simple—though most mentions are for breakfast/brunch.
+Figure 7: (a) Case 7; (b) Case 8